### v1

# Benchmarking Unified Face Attack Detection via Hierarchical Prompt Tuning

Abstract
Presentation Attack Detection (PAD) and Face Forgery Detection (FFD) are proposed to protect face data from physical media-based Presentation Attacks (PAs) and digital editing-based DeepFakes (DFs), respectively. However, isolated training of these two models significantly increases vulnerability towards unknown attacks, burdening deployment environments. The lack of a Unified Face Attack Detection (UAD) model to simultaneously handle attacks in these two categories is mainly attributed to two factors: (1) A benchmark that is sufficient for models to explore is lacking. Existing UAD datasets only contain limited attack types and samples, leading to the model’s confined ability to address abundant advanced threats. In light of these, through an explainable hierarchical way, we propose the most extensive and sophisticated collection of forgery techniques available to date, namely UniAttackDataPlus (UniAttackData+). Our UniAttackData+ encompasses $2,875$ identities and their $54$ kinds of corresponding falsified samples, in a total of $697,347$ videos. (2) The absence of a trustworthy classification criterion. Current methods endeavor to explore an arbitrary criterion within the same semantic space, which fails to exist when encountering diverse attacks. Thus, we present a novel Visual-Language Model-based Hierarchical Prompt Tuning Framework (HiPTune) that adaptively explores multiple classification criteria from different semantic spaces. Specifically, we construct a Visual Prompt Tree (VP-Tree) to explore various classification rules hierarchically. Then, by adaptively pruning the prompts, the model can select the most suitable prompts guiding the encoder to extract discriminative features at different levels in a coarse-to-fine manner. Finally, to help the model understand the classification criteria in visual space, we propose a Dynamically Prompt Integration (DPI) module to project the visual prompts to the text encoder to help obtain a more accurate semantics. Extensive experiments on 12 datasets have verified the potential to inspire further innovations in the UAD field.
Index Terms:
Unified Face Attack Detection, Face Anti-Spoofing, Face Forgery Detection, Prompt Tuning.
I Introduction
Face data are widely used in fields such as facial payment, mobile unlocking, and access authentication, due to the discriminative and easily obtainable facial biometric features. However, attackers can forge or steal identities in two ways to engage in illicit activities. As shown in the upper left of Fig. 2 (a), attackers can deceive face recognition systems by using physical media to impersonate others before information collection—a method known as physical attacks or presentation attacks. Additionally, attackers may employ digital forgery (commonly referred to as deepfake) to edit or manipulate collected images for fraudulent purposes. To address the aforementioned threats, Face Anti-Spoofing (FAS), also known as Presentation Attack Detection (PAD), and Face Forgery Detection (FFD) have been proposed to ensure the security of facial data. Essentially, both methods are formulated as binary classification tasks, where the model determines the authenticity of an image by outputting a probability indicating whether it is real or fake.
Researchers have proposed a wide variety of attack methods and have constructed multiple databases to advance the development of defense algorithms. Common physical presentation attacks are generally categorized into two types: 2D attacks (such as printed paper, fabric, screen replays) and 3D attacks (including 3D masks and silicone head models). Most models achieve authenticity classification by analyzing image texture details, focusing on inconsistent edges, screen moiré patterns, and other indicators. However, their generalization performance has been challenged. Consequently, domain generalization protocols such as ICMO and SCW have been proposed to evaluate model generalization. For instance, [1] utilized X, [2] leveraged X, and [3] employed X, while recent studies [4], [5], [6] have achieved notable results using vision-language models by employing X for X. Digital attack methods can be broadly classified into three categories: editing, adversarial attacks, and generative attacks. Digital editing attacks aim to modify image attributes (e.g., changing facial expressions, age) or perform face swapping (replacing the current face with another), video-driven manipulation (using a video to drive facial expressions), and more. Adversarial attacks can be divided into pixel-level and semantic-level attacks, which either modify or erase relevant pixels or alter high-dimensional semantics to manipulate the original facial data. Digital generative attacks, the most recent and innovative type, involve generating entirely new facial data using generative models. To combat digital attacks, researchers typically focus on detecting inconsistencies in pixel-level and semantic-level information to achieve robust classification performance.
Figure 1: The dashed lines represent the classification criteria that separate real and fake faces in the feature space. (a) Existing models are trained to draw a single ”line” in the feature space, leading to unpromising classification results. (b) Our HiPTune can adaptively integrate the classification criteria from different semantic spaces to achieve robust detection towards various attacks. Different colors represent distinct classification criteria.
Figure 2: (a) The Physical and Digital Attack construction approaches. Physical Presentation Attacks aim to steal identity information before the capture devices, and Digital Attacks attempt to edit the existing image to falsify the face data. (b) Our UniAttackData+ is constructed in an explainable hierarchical way. The coarse-to-fine categories provide researchers with prior knowledge which is helpful for the model designs. (b) A comparison of HiFAB with existing FAS, FFD, and UAD datasets shows that HiFAB surpasses all others in terms of attack types and video quantity. The radius of the circle represents the scale of the dataset, which is the number of videos contained in the dataset, corresponding to the horizontal axis of the figure. (Note: Grandfake is image-based) The vertical axis represents the number of attack types included in the dataset. (d) Sample demonstrations of the UniAttackData+. The samples are selected from the second level of the structure. The number following each category name indicates the number of attack types included.
Existing PAD and FFD models are typically trained and deployed separately, which inevitably reduces their robustness against unknown attacks and increases deployment complexity. When confronted with unknown attack samples or hybrid attack types, models trained independently exhibit poor robustness against such attacks. A unified model that can leverage the advantages of different attack types while maintaining a lightweight parameter size is urgently needed. We attribute the lack of such a unified model to two primary reasons:
(1) The lack of a comprehensive dataset that provides the model with a broader learning space. Although there are currently three unified face attack detection datasets (JFSFDB, Grandfake, UniAttack), they simply mix existing datasets that are tailored to specific attack types. This approach introduces substantial noise, such as variations in style, background, and identity, making it difficult to accurately evaluate the model’s superiority and generalization capability. Furthermore, as shown in Fig. 2 (b), their sizes are relatively small, containing limited attack types, and many of them are not publicly accessible, making them difficult to utilize effectively. To mitigate these issues, we previously proposed UniAttackData, a dataset where a total of 14 attacks (2 physical and 12 digital) are applied to the same face, creating an identity-consistent face attack dataset. However, its size and the diversity of attack types remain insufficient, resulting in models that tend to overfit the training set and perform poorly when encountering previously unseen attack types.
(2) The absence of a reliable classification criterion that can accurately categorize each sample. Existing methods typically aim to identify a single classification criterion within a single semantic space, as illustrated in Fig. 1 (a), where the model is trained to discover a single ’line’ that separates real and fake faces. However, in the context of UAD, this single classification criterion cannot robustly classify all sample types. From the perspective of data distribution, two major challenges arise: First, the diversity of attack types is constantly expanding, and these attacks increasingly resemble real samples across various dimensions. A single classification criterion is inadequate to address the rapidly evolving and sophisticated forgery techniques. Second, the significant feature differences between physical and digital attacks result in large intra-class variance among forged samples. The feature extractor struggles to effectively cluster them within a common space, making it difficult for a single classification criterion to determine the class of each sample accurately.
To address the dataset limitations mentioned above, we adopted the identity-consistent approach of UniAttackData and developed UniAttackDataPlus (UniAttackData+), which is currently the largest publicly available unified face attack detection dataset. It features the most extensive variety of attack types, minimal interference, and highly structured characteristics, making it a robust and comprehensive resource for training and evaluating face attack detection models. We collected data from $2,875$ participants representing three distinct ethnic groups (African, East Asian, and Central Asian), capturing $18,250$ authentic videos under various lighting conditions, backgrounds, and acquisition devices. For each participant, we applied $54$ different attack methods, including $14$ physical attacks and $40$ digital attacks, resulting in a total of $679,097$ forged videos, which is $32$ times more than the existing dataset. From each video, we randomly extracted at least $30$ frames to construct the dataset; samples can be found in Fig. 2 (c). The diverse acquisition backgrounds of authentic faces provide a more varied dataset, encouraging models to focus on discriminative features rather than irrelevant factors. Moreover, the extensive volume of data and the wide variety of attack types offer a more comprehensive and expansive learning space, motivating researchers to design more sophisticated and advanced models. Most importantly, UniAttackData+ applies a full range of attack methods to each unique identity, covering almost all known attack types. This not only prevents models from achieving high performance through ’identity memorization’ but also enhances the dataset’s applicability in real-world deployment scenarios. Additionally, our dataset is structured in a hierarchical manner, with all attack types categorized from coarse to fine, as depicted on the right side of Fig. 2 (a). This hierarchical structure provides researchers with a more interpretable framework for designing corresponding algorithms. For instance, our baseline method, HiPTune, leverages a similar hierarchical approach to develop a prompt-tuning method. Furthermore, we have designed four protocols to evaluate model performance under various scenarios, including unknown identities and unseen attack methods. We also introduced a cross-dataset testing benchmark to assess model generalization. Compared to previous datasets that simply merged existing datasets, UniAttackData+ offers a robust benchmark for the face data security community in a more interpretable hierarchical format, significantly promoting the development of more advanced algorithmic models.
To address the issue of insufficient robustness caused by a single classification criterion in unified attack detection, we propose a hierarchical prompt strategy named Hierarchical Prompt Tuning (HiPTune). HiPTune is designed to explore multiple classification criteria across different semantic spaces and adaptively integrate the most appropriate criterion for each sample. Specifically, following the hierarchical structure of our dataset (as shown on the right side of Fig. 2 (a)), we construct a Visual Prompt Tree (VP-Tree) as a carrier of multiple classification criteria. These learnable parameters, referred to as visual prompts, are concatenated with image tokens and fed into a fixed image encoder for training. These prompts serve as guiding symbols that direct the encoder to extract critical discriminative features, which we define as our ’classification criteria.’ The visual prompts are organized in a tree structure, progressing from coarse to fine, where each level represents a different granularity of classification criteria, and each leaf node corresponds to a specific criterion that is robust enough to distinguish certain fake samples from real faces. By inserting these visual prompts into different depths of the Transformer blocks in a top-down manner, we ensure that the model first learns to capture macro-level differences, achieving comprehensive detection of forgeries. As the feature extraction process deepens, finer-grained prompts focus on local details, enabling precise forgery discrimination. Shown as Fig. 1 (b), our HiPTune aims to utilize coarse-to-fine criteria to achieve a more adequate classification. This design endows the prompts with both global discriminative power and local specialization: top-level prompts enhance recall for unknown forgeries, while lower-level prompts improve fine-grained recognition of known categories. Moreover, to enrich the semantic information of these prompts, we introduce a Dynamic Prompt Integration (DPI) module, which employs a self-attention structure to adaptively integrate prompts from different levels and map them into the text embedding dimension. This cross-modal integration strategy broadens the semantic coverage of classification criteria, ensuring classification accuracy from multiple dimensions. Subsequently, we calculate the self-attention scores of the input samples, guiding the model to adaptively select the appropriate classification criteria at different levels for each sample. This adaptive strategy preserves model autonomy, avoiding excessive dependence on manual knowledge while uncovering potential connections between different attacks. It enables more robust detection tailored to each input sample.
To summarize, our contributions are as follows:
  • We create UniAttackData+, the largest publicly available unified face attack detection dataset, encompassing $54$ attacks per identity in an explainable hierarchical manner.
  • We proposed the Hierarchical Prompt Tuning (HiPTune) framework, which adaptively integrates classification criteria from different semantic spaces to establish a more robust criterion for each input sample.
  • We propose four intra-dataset protocols and nine cross-dataset benchmarks for the Unified Face Attack Detection community that promote the development of the detection model.
II Related Works
II-A Unified Face Attack Detection
II-A1 Datasets
Recently, several datasets, including Grandfake [1], JFSFDB [2], and UniAttack [3], have been introduced to address the challenges of UAD. These datasets combine existing PAD and FFD datasets to encompass a variety of attack types within a single dataset. Additionally, UniAttackData [4] was proposed as a UAD dataset with identity (ID) consistency, where each ID includes a complete range of attack types, thereby reducing ID and domain noise. The current UAD dataset not only lacks emerging generative forgery face samples but also outdated forgery algorithms corresponding to digitally edited faces, making it impossible to train a high-quality joint attack detection model. Our UniAttackData+ addresses this gap by incorporating a broader range of modern attacks and structuring them in an explainable manner, rather than merely aggregating datasets.
II-A2 Approaches
For the UAD task, Deb et al. [1] introduced a multi-task learning approach that treats each attack type as a separate detection task based on feature similarity. Yu et al. [2] focused on improving generalization by utilizing a multi-modal approach for joint task training. Fang et al. [4] leveraged the Vision-Language Model to capture both unified and specific knowledge, while Cao et al. [3] aimed to enhance face anti-spoofing by promoting the learning of compact, fundamental features of genuine faces. However, these methods rely on a single, generalized classification criterion, which often proves inadequate when faced with advanced attacks. In contrast, our HiPTune framework adaptively integrates the most suitable prompts from various semantic spaces, enabling more effective classification.
II-B Prompt Learning
PREFER [5] enhances prompt ensemble learning with a feedback mechanism for iterative refinement and prompt bagging, improving stability and performance across tasks. VPT [6] extends prompt learning from NLP to visual tasks. S-Prompt [7] clusters domain samples using K-NN, while C-Prompts [8] selects prompts from different domains for robust Domain Incremental Learning. CoOp [9] demonstrated that replacing manual prompts in CLIP with learnable contexts enhances performance via few-shot learning. CoCoOp [10] further adapts prompts to new domains with a Meta-Net, enhancing generalization.
However, existing methods overlook hierarchical prompt relationships, which capture both global and detailed knowledge. HiPro [11] organizes tasks hierarchically but lacks clarity between tasks. Our proposed HiPTune is able to adaptively select and integrate classification criteria from different semantic spaces by the VP-Tree, ensuring that the model not only learns to capture macro-level differences but also focuses on local details for the precise forgery discrimination.
Dataset Attack Type (each ID) # Datasets / Data # ID Physical Attacks Digital Attacks
Dataset Name No. # Categories No.
GrandFake Incomplete 6 sets: 789,412 (I) (Live: 341738, Fake: 447674) 96,817 SiW-M 128,112 (I) Adv (6) 116,641 (I)
DeepFake (6) 202,921 (I)
JFSFDB Incomplete 9 sets: 27,172 (V) (Live: 5,650, Fake: 21,522) 356 SiW 3,173 (V) DeepFake (4) 13,752 (V)
3DMAD 85 (V)
HKBU 588 (V)
MSU 210 (V)
3DMask 864 (V)
ROSE 2,850 (V)
UniAttackData Complete 1 set: 28,706 (V) (Live: 1,800, Fake: 26,906) 1,800 CASIA-SURF CeFA 5,400 (V) Adv (6) 10,706 (V)
DeepFake (6) 10,800 (V)
UniAttackData+ Complete 3 set: 697,347 (V) (Live: 18,250, Fake: 679,097) 2,875 CASIA-SURF 6,000 (V) Adv (16) 266,576 (V)
CeFA 9,000 (V) DeepFake (17) 242,859 (V)
HiFiMask 40,950 (V) Generation (7) 113,712 (V)
TABLE I: Comparison of existing UAD datasets. UniAttackData+ surpasses all other datasets in scale and diversity. ”V” and ”I” indicate video- and image-based counts, respectively. The numbers following each attack type in the ”# Catagories” row represent the number of algorithms included in that category.
Dataset # Attack Types # Videos
CASIA-SURF phy.6 21,000
CeFA phy.3 23,538
HiFiMask phy.5 54,600
SiW-v2 phy.14 1,700
MSU-MFSD phy.2 440
CASIA-FASD phy.3 600
ReplayAttack phy.2 1200
OULU-NPU phy.2 5940
FaceForensics++ digi.5 6,000
DeeperForensics-1.0 digi.1 60,000
DFDC digi.8 128,064
ForgeryNet digi.15 221,247
UniAttackData+ phy.10, digi.40 697,347
TABLE II: Comparison of existing PAD and FFD datasets. The scale of our UniAttackData+ is far more than the types in the existing PAD and FFD datasets. The ”#” represents ”the number of”.
III UniAttackDataPlus
III-A Advantages of UniAttackData+
Our UniAttackData+ dataset addresses five critical issues in current UAD datasets, i.e., insufficient sample size, limited attack coverage, and lack of interpretability, providing a comprehensive and valuable benchmark for the UAD community.
Scale. UniAttackData+ includes $2,875$ identities spanning three ethnicities, totaling $697,347$ video segments, with at least $25$ frames per video. Of these, $18,250$ are authentic face videos, 55,950 are physical attack videos, and $623,147$ are digital attack videos. Detailed data is presented in Tab. I and Tab. II. Our UniAttackData+ shows its superiority in separate PAD and FFD datasets, and also in the existing UAD datasets. The large dataset ensures a comprehensive and complete learning space, guaranteeing the effectiveness of detection models from a data perspective.
Attack Types. UniAttackData+ comprises $54$ attack types in total, including $14$ physical attacks and $40$ digital attacks, covering all known attack types in the field. This diversity ensures the model does not overfit to a specific attack type, thereby enhancing its generalization ability. All the attack methods included in our dataset were proposed after 2021, incorporating the latest digital generation attacks. UniAttackData+ ensures the data’s applicability in real-world scenarios, significantly expanding the range of attack types for model training and reducing vulnerability to emerging advanced attacks.
Interpretability. Existing UAD datasets merely merge different datasets without any correlation between the data, attack methods, and identities. This simplistic combination hinders the model from focusing on genuine forgery cues and greatly reduces model robustness in real-world applications. In contrast, UniAttackData+ preserves the identity consistency from UniAttackData while structuring all attack types hierarchically. Each attack method and sample is categorized from coarse to fine and labeled accordingly. This construction method makes full use of prior knowledge, offering researchers an actionable perspective to address UAD challenges. The hierarchical label collection and setup provide a more comprehensive experimental and analytical space, enabling deeper and more interpretable analyses of detection models and algorithms.
III-B Collection of UniAttackData+
Live Face source. The live faces in the UniAttackData+, enhancing diversity in terms of face identity, angle, expression, and ethnicity, are sourced from three public datasets: CASIA-SURF [12], which includes $1,000$ subjects; CASIA-SURF CeFA [13] (CeFA), comprising 1,800 subjects from three ethnicities - African, East Asian, and Central Asian, with $600$ subjects each ethnicity; and CASIA-SURF HiFiMask [14] (HiFiMask), which contains $75$ subjects. These varied sources significantly enhance the dataset’s diversity across multiple dimensions. We merge three existing datasets for the live face data to obtain $2,875$ distinct identities. For CASIA-SURF, the Intel RealSense SR300 camera captures collectors turning left or right, moving up or down, and approaching or moving away from the camera. The same collection method is used for the CASIA-SURF CeFA dataset. HiFiMask contains multiple lighting conditions, such as white, green, and periodic three-colors. Additionally, supplementary light sources from six directions are used to simulate real-world application scenarios.
For dataset preprocessing, we apply MTCNN [15] for face detection and cropping on the real face data from the original datasets, with the scale parameter set to 1.3. We then manually assign class labels, attack type labels, and hierarchical labels to facilitate research.
III-C Spoofing Approach
To ensure diversity in physical attack methods within the UniAttackData+ dataset, we incorporate $14$ physical face spoofing techniques, including $11$ 2D presentation attacks and $3$ 3D mask attacks. The 2D attacks consist of $6$ paper-cutout variations and $2$ print attack types, as well as $3$ video-replay attacks to enrich the dataset diversity further. For 3D mask attacks, UniAttackData+ includes high-fidelity masks made from three different materials: transparent, plaster, and resin. These masks are recorded under varied lighting conditions and with multiple imaging devices for added variation.
2D Attacks. We include two kinds of print attacks recorded in the outdoor and indoor environments. Moreover, three kinds of screen replay attacks from different devices are utilized. For the cutouts, the printed flat or curved face images will be cut into eyes, nose, mouth areas, or their combinations, generating 6 different attack ways. Specifically, the six ways are: (1) One person holds his/her flat face photo where eye regions are cut; (2) One person holds his/her curved face photo where eye regions are cut; (3) One person holds his/her flat face photo where eye and nose regions are cut; (4) One person holds his/her curved face photo where eye and nose regions are cut; (5) One person holds his/her flat face photo where eye, nose, and mouth regions are cut; (6) One person his/her curved face photo where eye, nose, and mouth regions are cut.
3D Attacks. For the 3D attacks, we construct three kinds of masks with different materials: transparent, plaster, and resin. All masks are worn on a real person’s face and uniformly dressed to avoid the algorithm looking for clues outside the head area.
Category Type Methods # Videos Total
2D Print Cloth-Indoor 1,800 3,600 55,950
Cloth-Outdoor 1,800
Replay pic-phone 1,800 5,400
screen-phone 1,800
screen-rs 1,800
Cutouts Cutout1 1,000 6,000
Cutout2 1,000
Cutout3 1,000
Cutout4 1,000
Cutout5 1,000
Cutout6 1,000
3D Transparent 13,650 40,950
Plaster 13,650
Resin 13,650
TABLE III: The detailed statistics of physical attacks in UniAttackData+. The ”#” represents ”the number of”.
III-D Forgery Approach
III-D1 Digital Manipulation Attacks
Digital Manipulation Attacks in UniAttackData+ encompasses $17$ specific methods grouped into three types of attacks: (1) Face Swapping [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], where identity features of a source face, $I_{\text{source}}$, are transferred into a target face, $I_{\text{target}}$, producing a composite image, $I_{\text{composite}}=f(I_{\text{source}},I_{\text{target}})$. (2) Facial Attribute Editing [27, 28], which alters specific attributes of a face, such as age, expression, or gender, expressed as $I^{\prime}=g(I,A_{\text{target}})$, where $A_{\text{target}}$ denotes the desired attribute. (3) Video-Driven Facial Animation [29, 30], which generates facial movements in static images using motion parameters $M$ from a driving source, represented by $I_{\text{animated}}=h(I,M)$.
Category Type Methods # Videos Total
Digital Manipulation Attribute-edit SAFA [27] 15,998 32,659 242,859
StarGAN [28] 16,661
Face-Swap FaceDancer [16] 16,168 176,878
GHOST [17] 13,135
SimSwap [18] 10,563
InsightFace [19] 13,264
BlendFace [20] 9,645
DeepLiveCam [31] 12,059
REFace [21] 12,756
FOMM [22] 16,661
InfoSwap [23] 16,534
MegaFS [24] 14,334
HifiFace [25] 16,661
DiffSwap [26] 13,761
Roop [32] 11,337
Video-Driven DaGAN [29] 16,661 33,322
OneShotTH [30] 16,661
TABLE IV: The detailed statistics of digital manipulation attacks.
Face-Swap. FaceDancer [16] utilizes Adaptive Feature Fusion Attention (AFFA) and Interpreted Feature Similarity Regularization (IFSR) to achieve seamless face-swapping with high-fidelity identity transfer and pose preservation. GHOST [17] introduces architectural enhancements like an eye-based loss function, face mask smoothing, and stabilization techniques to achieve high-fidelity image-to-image and image-to-video face swapping with improved identity preservation and gaze alignment. SimSwap [18] employs an ID Injection Module (IIM) for feature-level identity transfer and a Weak Feature Matching Loss to preserve target facial attributes, enabling high-fidelity and generalized face swapping. BlendFace [20] introduces a novel identity encoder trained on blended images to mitigate attribute biases, enabling better identity-attribute disentanglement for improved face-swapping while maintaining competitive performance. Deep-Live-Cam [31] provides tools for animating custom characters and modeling tasks, with built-in safeguards to prevent processing of inappropriate content, emphasizing ethical and lawful use, user responsibility, and adherence to legal requirements, including potential watermarking or project suspension if necessary. REFace [21] leverages diffusion models to frame face-swapping as a self-supervised inpainting task, incorporating DDIM sampling, CLIP feature disentanglement, and mask shuffling for enhanced identity transfer, realism, and versatility, including head and accessory swapping. FOMM [22] introduces unsupervised motion representations for animating articulated objects by identifying and tracking distinct parts, using region-based descriptions to disentangle shape and pose, and modeling global motion with affine transformations for improved animation fidelity. InfoSwap [23] uses an information bottleneck and identity contrastive loss to enhance identity disentanglement, generating high-quality, identity-discriminative swapped faces. MegaFS [24] introduces a Hierarchical Representation Face Encoder (HieRFE) and a Face Transfer Module (FTM) for high-quality megapixel-level face swapping, leveraging StyleGAN2 for stable synthesis and low memory usage. HifiFace [25] employs 3D shape-aware identity with geometric supervision and a Semantic Facial Fusion module for adaptive blending, achieving superior identity preservation and photo-realistic face-swapping results. DiffSwap [26] reframes face swapping as a conditional inpainting task using diffusion models, introducing a midpoint estimation method for efficient training and leveraging 3D-aware landmarks for controllable, high-fidelity, and shape-preserving face swapping. Roop [32] takes a video and replaces the face in it with a face of your choice. You only need one image of the desired face.
Attribute-edit. SAFA [27] employs 3D morphable models (3DMM) for facial structure modeling and inpainting techniques for occlusion recovery, enabling realistic and structure-aware face animation even for significantly deviated poses. StarGAN [28] employs a unified model architecture that enables multi-domain image-to-image translation within a single network, achieving superior scalability, robustness, and image quality compared to existing models.
Video-Driven. DaGAN [29] leverages self-supervised depth learning to recover dense 3D facial geometry, using it for keypoint estimation and 3D-aware cross-modal attention to guide motion field generation, enabling realistic and accurate talking head video synthesis. OneShotTH [30] uses a novel keypoint representation to decompose identity and motion information unsupervisedly, enabling high-quality talking-head video synthesis with reduced bandwidth and interactive head rotation for enhanced video conferencing experiences.
III-D2 Digital Adversarial Attacks
UniAttackData+ includes $16$ attacks designed to mislead machine learning models through subtle, often imperceptible perturbations. These attacks are categorized into pixel-level attacks [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44], which modify raw image pixels to create adversarial examples, and semantic-level attacks [45, 46, 47, 48], which alter high-level features to change model interpretations. Pixel-level attacks rely on precise, gradient-based adjustments, while semantic-level attacks manipulate intermediate representations or latent spaces, increasing their transferability and effectiveness.
Category Type Methods # Videos Total
Digital Adversarial Pixel_level ALMA [33] 16,661 183,271 266,576
AdvDrop [34] 16,661
AI_FGTM [35] 16,661
ILA-DA [36] 16,661
AutoAttack [37] 16,661
PGD [38] 16,661
Jitter [40] 16,661
NIFGSM [41] 16,661
PIFGSMpp [42] 16,661
Pixel [43] 16,661
VMIFGSM [44] 16,661
Semantic-Level Demiguise [45] 16,661 83,305
SSAH [46] 16,661
DeCoWA [39] 16,661
PNA_PatchOut [47] 16,661
TGR [48] 16,661
TABLE V: The detailed statistics of digital adversarial attacks.
Pixel-Level. ALMA [33] utilizes Augmented Lagrangian-based attack for efficient, minimally perturbed adversarial examples, combining generality and speed. AdvDrop [34] introduces a novel adversarial attack by dropping imperceptible details from images instead of adding disturbances, challenging DNN robustness and evading existing defense mechanisms. AI-FGTM [35] improves adversarial example generation with dynamic step size, smaller kernels, and a tanh-based gradient method, achieving higher transferability, indistinguishability, and attack success rates than state-of-the-art methods. ILA-DA [36] enhances adversarial example transferability by combining automated image transformations, reverse adversarial updates, and attack interpolation, significantly outperforming prior methods in both defended and undefended scenarios. AutoAttack [37] introduces an ensemble of parameter-free adversarial attacks, improving PGD with optimized step size and objective function to rigorously evaluate adversarial robustness, uncovering weaknesses in many published defenses. PGD [38] explores adversarial robustness through robust optimization, introducing methods for training and attacking neural networks with universal security guarantees against first-order adversaries, significantly enhancing resistance to adversarial attacks. Jitter [40] proposes a novel loss function that improves attack efficiency and success rates by better exploring the perturbation space and addressing model confidence issues. NI-FGSM [41] leverages Nesterov accelerated gradients, and SIM, optimizing perturbations across scaled copies, to enhance the transferability of adversarial examples, achieving higher success rates on defense models. PIM++ [42] enhances adversarial transferability by using a patch-wise iterative method with an amplified step size, a projection kernel for gradient redistribution, and temperature adjustments to balance transferability and white-box attack performance. Pixel [43] introduces a black-box attack that rearranges a small number of pixels to achieve high attack success rates across models and datasets, requiring few iterations and minimal perceptual changes. VMI-FGSM [44] enhances gradient-based adversarial attacks by stabilizing update directions with gradient variance, significantly improving transferability.
Semantic-Level. Demiguise Attack [45] crafts unrestricted, photorealistic perturbations based on perceptual similarity, improving fooling rates, transferability, and robustness against defenses while maintaining human imperceptibility and simulating real-world conditions. SSAH [46] attacks semantic similarity in feature representations for cross-dataset adversarial generalization and employs a low-frequency constraint to generate imperceptible perturbations, achieving high transferability and perceptual similarity. DeCoWA [39] introduces Deformation-Constrained Warping (DeCoW) to enhance adversarial transferability across model genera by augmenting inputs with controlled elastic deformations, effectively attacking both CNNs and Transformers across various tasks. TGR [48] reduces gradient variance in Vision Transformer (ViT) blocks token-wise, generating more transferable adversarial samples and improving attack performance. PNA-PatchOut [47] combines Pay No Attention (PNA), skipping attention gradients, and PatchOut, optimizing random patch subsets, to enhance adversarial transferability across Vision Transformers (ViTs) and CNNs.
III-D3 Digital Generation Attacks
UniAttackData+ comprise $7$ types leveraging diffusion models to produce highly realistic synthetic content. These attacks are divided into three categories: Identity-Consistent Generation, Style Transfer, and Prompt-Driven Generation. Identity-Consistent Generation [49, 50, 51] seeks to maintain a subject’s identity across various contexts, modeled as $I_{\text{gen}}=G(I_{\text{id}},z)$, where $I_{\text{id}}$ is the identity input and $z$ represents variations such as pose or expression. Style Transfer, as in PhotoMaker [52], adapts identity features to different styles, represented as $I_{\text{styled}}=G(I_{\text{id}},\mathbf{S}_{\text{style}})$, to create visually diverse outputs. Prompt-Driven Generation [53] utilizes textual or image prompts to guide synthesis, expressed as $I_{\text{prompted}}=G(P_{\text{text}},P_{\text{img}})$.
ID-Consistant. ConsistentID [49] enhances personalized portrait generation with a multimodal facial prompt generator for fine-grained control and an ID-preservation network with facial attention localization, achieving superior ID consistency and precision from a single reference image. PuLID [50] introduces a tuning-free ID customization method for text-to-image generation, combining a Lightning T2I branch with contrastive alignment and ID losses to ensure high ID fidelity and minimal disruption to original image elements. InstantID [51] introduces a plug-and-play module with IdentityNet, leveraging strong semantic and weak spatial conditions to achieve high-fidelity, single-image-based identity personalization, seamlessly integrating with pre-trained diffusion models like SD1.5 and SDXL.
Style-Transfer. PhotoMaker [52] introduces a stacked ID embedding to encode multiple ID images into a unified representation, enabling efficient personalized text-to-image generation with high ID fidelity, text controllability, and speed, supported by an ID-oriented data construction pipeline.
Prompt-Based. IP-Adapter [53] introduces a decoupled cross-attention mechanism to enable lightweight image prompting for text-to-image diffusion models, achieving multimodal generation with only 22M parameters, while maintaining compatibility with text prompts, structural controls, and custom models.
Category Type Methods # Videos Total
Digital Generation ID-Consistent ConsistentID [49] 15,436 46,321 113,712
PuLID [50] 16,210
InstantID [51] 14,675
Style Transfer PhotoMaker [52] 18,192 18,192
Prompt Based IP-Adapter_kolors [53] 16,449 49,199
IP-Adapter_sdplusface [53] 16,449
IP-Adapter_sdpxlface [53] 16,301
TABLE VI: The detailed statistics of digital generation attacks.
III-E Protocol
ID Partition Protocol $P1$. Simply merging existing PAD and FFD datasets can cause models to focus on ID noise rather than discriminative features. To address this, we leverage the ID consistency of UniAttackData+. In this protocol, we train the model on a selected subset of identities, using the remaining identities for validation and testing. Each ID contains all attack types.
Cross-attack methods Protocol $P2$. In real-world deployment scenarios, a myriad of attack methods may emerge. Therefore, models should not only fit to seen attack methods but must also be capable of handling unseen attack methods. This protocol is designed to evaluate the model’s generalization ability across different attack methods. We adopted all identities for training and randomly selected 50% of attack methods as the training set, 20% as the validation set, and 30% as the test set. For instance, in the case of digital attacks, we used the first seven types of digital editing face swap attacks as the training set, two types for validation, and four types for testing, in order to assess the model’s performance on unseen face swap attacks.
Cross-attack types Protocol $P3$. In many cases, models are trained only on simpler attack types but are tested against more advanced and complex attack methods. This protocol aims to assess the model’s generalization ability from simpler to more sophisticated attacks. We trained the model using all identities, with 70% of physical 2D attacks, digital editing, and adversarial attacks from each identity as the training set, and 30% as the validation set. Physical 3D attacks and digital generative attacks were used as the test set, to evaluate the model’s generalization ability on unseen, more advanced attack types.
IV Hierarchical Prompt Tuning (HiPTune)
Figure 3: (a) The VP-Tree establishes a hierarchical prompt tree for all attack types, organized into three levels of grouping categories. It assigns a learnable prompt for each attack type and integrates the hierarchical prompts into the final output. (b) The APP employs a hierarchical prompt pruning mechanism to select the most promising prompt for each sample. (c) The DPI aims to aggregate all candidate prompts from different levels of VP-Tree and projects them into the text encoder for subsequent similarity computation.
IV-A Preliminaries
Vision-Language Models (VLM) like CLIP [54] utilize both text encoder $\mathcal{T}(\cdot)$ (like Transformer [55]) and image encoder $\mathcal{I}(\cdot)$ (like ViT [56]), where the input is the (text, image) pair, to leverages broader supervision. The outputs of both encoders are then dot-multiplied to obtain the probability score $p$:
$p(y=i\mid\bm{x})=\frac{\exp\left(\cos\left(\bm{w}_{\bm{i}},\bm{f}\right)/\tau\right)}{\sum_{j=1}^{K}\exp\left(\cos\left(\bm{w}_{\bm{j}},\bm{f}\right)/\tau\right)},$ (1)
where $\left\{\bm{w}_{i}\right\}_{i=1}^{K}$ denotes the set of weight vectors generated by the text encoder, and $\bm{f}$ represents the features extracted by the image encoder. $K$ is the number of classes, and $\tau$ refers to the temperature parameter learned by CLIP, $\cos(\cdot,\cdot)$ means the cosine similarity. Moreover, CoOp [57] explores prompt engineering in the VLM, constructing learnable prompts $\bm{t}=[\mathrm{V}]_{1}[\mathrm{~{}V}]_{2}\ldots[\mathrm{~{}V}]_{M}[\mathrm{CLASS}]$ and feeding to the fixed Text Encoder. The $[\mathrm{V}]_{m}(m\in\{1,\ldots,M\})$ is the vectors with the same dimension as word embeddings, and $M$ is a hyperparameter specifying the number of context tokens. The prediction probability is further computed as follows:
$p(y=i\mid\bm{x})=\frac{\exp\left(\cos\left(\mathcal{T}\left(\bm{t}_{i}\right),\bm{f}\right)/\tau\right)}{\sum_{j=1}^{K}\exp\left(\cos\left(\mathcal{T}\left(\bm{t}_{j}\right),\bm{f}\right)/\tau\right)}.$ (2)
Visual Prompt Tuning (VPT) aims to fine-tune the pre-trained models with only a few learnable parameters. Given the input image tokens $\bm{x}_{i},i\in{1,...,l}$ with length $l$, the trainable vectors $\bm{P}=\left\{\bm{p}^{k}\in\mathbb{R}^{d}\mid k\in\mathbb{N},1\leq k\leq p\right\}$, where d is the dimension of the patch embeddings and $p$ is the number of prompts, are cat with the original tokens and fed into the fixed encoders $Input=cat[\bm{x}\cdot\bm{P}]$.
IV-B Overview of HiPTune
We proposed a HiPTune that adopts hierarchical prompts to find multiple classification criteria from different semantic spaces and further integrates them, aiming to form more robust classification criteria for each input sample. Shown in Fig. 3, first, we construct the Visual Prompt Tree to gain the various semantics spaces tailored for each attack detection task. Then, given the input image $\bm{x}$, we adaptively select the most suitable prompt from each level of the VP-Tree by our Adaptive Prompt Pruning module. In addition, we integrate the prompt candidates from each selected level of the APP to describe the complete features of the input sample and feed it to the text encoder for better alignment.
IV-C Visual Prompt Tree (VP-Tree)
Define the levels. We can obtain a hierarchical relationship between each attack type based on the spoof and forgery types. Apart from the subjective categorization from the dataset, the features of different attacks demonstrate a more compact grouping relative to other categories while maintaining distinct independence within each category. For a VP-Tree with $N$ degrees, the level ${Tree}^{i}(i\in{1,2,...,N})$ has $n$ prompts, and $n$ may vary according to the set-ups. In our case, we define $3$-levels for the VP-Tree shown as Fig. 3 (a). The first level distributes the most obvious data as live, physical, and digital attack subtrees. The physical subtree is first separated into 2D and 3D attacks, which represents the second level of the VP-Tree. Print, replay, and cutouts are three key types for the level 3 prompts of a 2D attack. As for the 3D attacks, the masks’ material is always considered to be the spoof clues, so we define transparent, plaster, and resin prompts within the 3D attack prompt. Concerning the digital subtree, we categorize the digital attacks into $3$ kinds according to their forgery techniques. The digital manipulation method includes attribute-editing, face-swap, and video-driven faces. Adversarial attacks are divided into pixel-level and semantic-level attacks for forgery areas. We also define $3$ sub-prompts for the digital generation attacks: ID-consistent generation, style transfer, and prompt-based editing forgeries.
Initialize the prompts. After we obtain the VP-Tree’s structure, we define the prompts for each level with random initialization in the following form:
$\bm{p}=[\mathrm{V}]_{1}[\mathrm{~{}V}]_{2}\ldots[\mathrm{~{}V}]_{l},$ (3)
where each $[\mathrm{V}]_{l}(l\in\{1,\ldots,L\})$ is the learnable vectors with the length of $L$ at the same dimension with the patch embeddings (i.e., 768 for CLIP image encoder).
Train and integrate the prompts. In order to force each prompt to learn the classification criteria specific to each attack type, we train the prompts in a supervised manner during the first stage. Given the input $\bm{x}\in\mathbb{R}^{C\times H\times W}$, where $C,H,W$ represent the channel, height, and width of the input images, we manually record the attack types and select the corresponding prompts at each level accordingly. For the live face images, since we train our model with (live, fake) pairs, we choose the same prompt as the partner. Then, we average the selected prompts to form the final prompt and link with the image tokens. The final input of the image encoder is formulated as Eq. 4:
${\mathcal{I}}_{input}=cat[(\bm{x}_{cls},\bm{x}_{1},...,\bm{x}_{l})\cdot\bm{p}],$ (4)
where $(\bm{x}_{cls},\bm{x}_{1},...,\bm{x}_{l})\in\mathbb{R}^{l+1,dim}$ is the image tokens after the patch embeddings. $l$ represents the length of image tokens, and $dim=768$ is the embedding dimension. The $\bm{p}$ is the final output of the integrated prompts as Eq. 5
$\bm{p}=mean(\bm{p}_{1}+\bm{p}_{2}+\bm{p}_{3}),$ (5)
where $\bm{p}_{1},\bm{p}_{2},\bm{p}_{3}$ are the selected prompts at level $1,2,3$ and $mean(\cdot)$ represents the average of the inputs.
After fully training the VP-Tree to represent the classification criteria for multiple semantic spaces, the Adaptive Prompt Pruning (APP) module will learn to select the most suitable prompts for each sample.
IV-D Adaptive Prompt Pruning (APP)
As shown Fig. 3 (b), aiming to explore classification criteria from different semantic spaces, the APP module adopts a hierarchical pruning mechanism to select the most suitable prompt for each sample. Given the input image $\bm{x}\in\mathbb{R}^{C\times H\times W}$, the patch embedding first projects the image into the visual embedding dimension and reshapes $\bm{x}$ into a series of image token with length $L$: $\bm{x}:\mathbb{R}^{C\times H\times W}\rightarrow\mathbb{R}^{L\times D}$. In our case, $C=3$ represents the channels for the image, and $H,W$ are both $244$ for the input image, the token’s length $L=196$, and the dimension of visual embedding is $D=768$. After adding the positional embeddings, a class token is linked with the image token, and then the image token $\bm{x}\in\mathbb{R}^{197\times 768}$ is fed to the APP.
Level-1 prompt pruning. Given the significant distribution variance between physical and digital attacks, it is simple to determine whether physical or digital criteria should be used to classify this sample. Thus, we utilize a simple linear layer to select the level 1 prompt:
$G_{1}(\bm{x})=\mathbf{W}_{1}\cdot\bm{x}+bias_{1},$ (6)
where $G_{1}$ is the gate function of the level-1 prompt pruning and $\mathbf{W}_{1},bias_{1}$ are learnable weights and bias. The level-1 router first projects each patch’s features into three alternatives using $G_{1}$ and then calculates the mean value over all tokens. After applying Softmax between three classes, we choose the $TOP_{K=1}$ candidate prompt. The whole procedure can be described as:
$i^{p}_{1}=TOP_{k=1}(softmax(mean(G_{1}(\bm{x}),197))),$ (7)
where $i^{p}_{1}$ denotes the index of the selected prompt, $mean(\cdot,\cdot)$ represents the mean calculus over specific dimension. Furthermore, the selected prompt is linked with the original image tokens and forwarded to the next level.
Level-2 prompt pruning. Regarding the level-2 prompt pruning, it is challenging to separate samples from attack types within each category if we only rely on fixed visual embeddings. In light of this, we apply Central Difference Convolution (CDC) [58] and the Frequency-Adaptive Dilated Convolution (FADC) [59] for level-2 prompt pruning in the physical and digital branches, respectively.
After the CDC or the FADC, the output feature map is sent to a linear layer, and then the tokens’ scores are averaged to select the candidate prompt further. The level-2 gate functions can be expressed as:
$G_{2}^{0}$ $=\mathbf{W}_{2}^{0}\cdot CDC_{2}^{0}(\bm{x})+bias_{2}^{0}$ (8)
$G_{2}^{1}$ $=\mathbf{W}_{2}^{1}\cdot FADC_{2}^{1}(\bm{x})+bias_{2}^{1},$
and the selected prompt can be denoted as:
$i^{p}_{2}=TOP_{k=1}(softmax(mean(G_{2}(\bm{x}),197))).$ (9)
The level-1 and level-2 prompts are combined averagely and concat with the image tokens and fed to the next router.
Level-3 prompt pruning. The level-3 prompt pruning procedure is the same as the level-2 pruning process. The gate functions can be described as follows:
$G_{3}^{i}=\mathbf{W}_{3}^{i}\cdot Conv_{3}^{i}(\bm{x})+bias_{3}^{i},$ (10)
where $i\in{1,2,3}$ for physical 2D/3D, digital manipulation, generation, and $i\in{1,2}$ for digital adversarial attacks.
Finally, the selected prompts from all three levels are averaged together and concated with the original image tokens. For the live images, the live prompt from the first level will be directly concated with the tokens and fed to the image encoder.
IV-E Dynamic Prompt Interaction (DPI)
CLIP’s default templates,  e.g., “This is a photo of [CLASS]”, fail to describe the detailed semantic information for the multiple attack types, which caused the performance drops encountering advanced attacks. Moreover, the unified prompts in the CoOp still can not represent the differences between various attacks and the live faces. However, as mentioned above, our VP-Tree constructs a hierarchical prompt tree for the different classification criteria, and our APP obtains the probabilities of the suitable candidate prompts. Thus, we proposed a Dynamic Prompt Interaction (DPI) to project the visual prompts to the text encoder to fully describe the abundant features corresponding to each category.
Specifically, first, we directly pass the live prompts from the first level as the live class text prompts. As for the fake prompts, we apply the probabilities that the app has predicted to the candidate prompts and combine them weightedly. For the first level, we apply the softmax function only to physical and digital prompt scores to obtain the binary probabilities of two categories. Then, we weighted sum the prompts to get a weighted level-1 fake prompt. In the second level, due to the router from the first level choosing a subtree for the sample and needing a possible fake prompt specific to the input sample, we repeat the weighted sum of candidate prompts until the router has not predicted the further probabilities. Lastly, we form prompts for both live and fake faces and send them to the text encoder for further similarity computation.
IV-F Objectives and Inference
The training objective for the first training stage is to help the prompts to learn the semantic information and classification criteria specific to the attack type. The second stage aims to train the APP to select the most suitable prompts tailored for each sample successfully. The loss for both stages is:
$\mathcal{L}=\mathcal{L}_{CE}+\mathcal{L}_{triple},$ (11)
where $\mathcal{L}_{CE}$ represents the cross-entropy loss and $\mathcal{L}_{triple}$ is the Asymmetric Triplet Loss [60]. The reason for applying Asymmetric Triplet Loss is to separate the different prompts to preserve their unique classification criteria while combat the relatively similar prompts.
V Experiments
V-A Experiment Settings
Datasets, Protocols, and Evaluation Metrics. We evaluated the effectiveness of the proposed UniAttackData+ dataset and the robustness of the HiPTune algorithm using three existing UAD datasets. For JFSFDB, we conducted both internal and external testing within its original FAS and FFD datasets, as well as within the joint training dataset. For UniAttackData, we tested using its original p1, p2.1, and p2.2 subsets. For our proposed UniAttackData+, we performed training and testing based on the four protocols we introduced. Additionally, we established a UAD-Benchmark to assess the model’s performance in cross-dataset testing. We evaluate model performance with four metrics: (1) ACER (average of FRR and FAR), (2) AUC (theoretical model performance), (3) EER (optimal threshold), and (4) ACC (overall accuracy).
Implementation Details. We use a $3$-level VP-Tree with a prompt length of $60$, a $\theta$-value of CDC set to $0.7$, and a fixed CLIP model with ViT/b-16 backbone, classifying ”live” and ”fake.” The initial learning rate is $10^{-5}$ with the Adam optimizer. Images are resized to $224\times 224$ with random cropping and flipping. VP-Tree training runs on an NVIDIA GeForce RTX 4090 for $100$ epochs (batch size: $32$), followed by full model training for $200$ epochs (batch size: $32$).
V-B Results on UniAttackData+
p1-ID p2-cross_attack_methods p3.1-cross_attacks(low2high) p3.2-cross_attacks(high2low)
Model ACER AUC ACC EER ACER AUC ACC EER ACER AUC ACC EER ACER AUC ACC EER
CLIP-V 9.36 92.46 88.05 9.36 28.77 71.99 78.49 28.77 49.87 43.12 58.08 49.87 51.17 47.35 86.76 51.17
CLIP 9.93 95.22 83.75 9.93 22.89 72.79 76.15 22.89 34.37 58.09 59.9 34.37 53.05 46.95 82.68 53.05
CoOp(unified) 14.76 90.89 83.92 14.76 43.14 59.64 47.83 43.14 72 20.89 27.18 72 45.06 53.7 59.2 45.06
CoOp(specific) 15.26 90.8 84.48 15.26 44.73 55.44 48.5 44.73 71.57 21.76 24.79 71.57 42.82 57.26 57.83 42.82
CDCN++ 12.77 94.53 90.89 11.72 15.23 91.78 84.72 15.22 37.55 67.9 58.67 38.46 42.09 62.16 70.55 39.66
CFPL-FAS 13.93 93.77 96.235 11.65 16.1 86.34 83.09 16.1 29.6 73.5 71.46 29.6 53.2 48 84.98 50.21
MoE-FFD 17.3 91.9 96.5 15.69 30.53 70.09 92.82 31.39 46.01 51.27 89.14 47.97 54.85 45.11 88.81 54.85
LLA-Net 10.57 95.64 95.97 7.61 18.34 82.97 89.86 21.31 36.58 57.02 68.25 42.2 54.04 45.94 87.71 54.04
FA-CLIP 8.26 93.96 94.53 8.45 35.81 63.89 88.51 47.37 44.44 48.34 56.95 41.05 51.22 48.78 90.38 99.64
HiPTune 7.31 97.87 90.02 7.61 14.26 92.5 90.26 14.71 25.68 77.62 69.34 26.02 40.13 65.89 92.66 40.18
TABLE VII: Results on UniAttackData+. CoOp(unified/specific) represents the CoOp without/with the class-specific context.
We employed a range of CLIP-based baseline models (including CLIP-V, CLIP, CoOp without class-specific context, and CoOp with class-specific context) to evaluate the performance of shared-space learning models on our dataset. In addition, we assessed algorithms specifically designed for PAD/FFD, such as CFPL-FAS, MoE-FFD, and LLA-Net, to determine their generalization capabilities. We also compared our approach with the recently released $FA^{3}-CLIP$, an algorithm specifically designed for unified attack detection. The detailed experimental results are presented in Table VII.
Authenticity-Only Testing Protocol $p1$:
This protocol primarily aims to evaluate the model’s true capability in distinguishing between real and fake samples. As shown in the experimental results table (Table VII), models that attempt to update the encoder generally achieve better performance. This is because, within UniAttackData+, a shared feature space still exists where different attack types can be clustered, and enhancing the image encoder’s feature extraction capability improves model performance. However, these models have a large number of parameters, making them unsuitable for deployment in real-world environments.
Conversely, CoOp, which relies on prompt matching, attempts to map fixed features to relevant regions. However, given the vast and unpredictable variety of attack types, a single prompt is insufficient to address the diversity and sophistication of these attacks. Models specifically designed for PAD and FFD focus solely on the manipulated regions of the respective images, neglecting the common forgery cues and underlying processes shared across UAD tasks. Notably, CDCN++, a purely convolutional model, also demonstrates decent performance, highlighting the importance of local region analysis in UAD tasks. This suggests that local feature extraction remains crucial for robust attack detection.
Our HiPTune adopts a visual prompt approach, which not only guides the encoder in feature extraction but also maintains model efficiency. Furthermore, our hierarchical prompt tree encapsulates multi-level, fine-grained classification criteria, achieving highly precise classification performance.
Real-World Robustness Testing Protocol $p2$:
This protocol is widely applicable in real-world scenarios, aiming to evaluate the model’s performance on seen attack types (such as physical 2D attacks) and unseen attack methods (such as face-swapping techniques like Roop and SimSwap). Similar to protocol p1, models that train the encoder generally achieve superior performance. This is because identifying a shared feature space helps the model generate a general mapping for previously unseen methods, effectively leveraging existing knowledge to handle new attack techniques.
In contrast, models like CoOp, which rely on prompt matching, are more prone to overfitting the training data. As a result, their detection performance significantly declines when encountering new data distributions. For instance, CDCN++, a model specifically designed for PAD, utilizes a central difference convolutional structure that emphasizes local detail information, achieving the best results in our benchmark experiments. This finding highlights the importance of focusing on local detail information to ensure model generalization in UAD tasks.
Our HiPTune model adopts a visual prompt approach, introducing prompts on the image encoder side to guide the frozen encoder in extracting relevant features while maintaining efficiency. Unlike conventional prompt-matching methods, our prompts serve only as guidance, avoiding excessive dependence on existing knowledge for decision-making. This design not only enhances the model’s generalization capability against unseen attack methods but also ensures robustness and interpretability.
Upward Compatibility Protocol $p3.1$:
This protocol aims to assess whether models trained on simple attack types can maintain high robustness in testing environments featuring complex and advanced attacks. We observe a sharp performance decline in both CoOp models, which we attribute to the absence of encoder guidance. Relying solely on text-side mapping relationships fails to achieve strong reasoning and generalization capabilities.
In contrast, our HiPTune model, which adaptively integrates different types of prompts, demonstrates superior performance. Despite being trained on limited samples, HiPTune can make judgments based on existing higher-level prompts while leveraging data from other attack types to assist in inference. This design ensures that the model retains strong generalization even in the presence of advanced attack types.
Downward Compatibility Protocol $p3.2$:
This protocol aims to evaluate whether models can leverage advanced knowledge to accurately detect simpler attack types. As shown in the table, most models exhibit poor performance, indicating that existing methods are insufficient to effectively utilize anti-forgery knowledge derived from advanced digital generative attacks. This also reveals the potential non-transferability between these two types of knowledge.
In contrast, our HiPTune model demonstrates superior performance. It effectively borrows knowledge from higher-level prompts and dynamically combines existing classification criteria, highlighting its advanced design. This adaptive approach allows HiPTune to generalize knowledge from complex attacks to simpler ones, ensuring robust performance across diverse scenarios.
V-C Results on UAD-Benchmark
JFSFDB-UniAttackData+ UniAttackData-UniAttackData+ JFSFDB+UniAttackData-UniAttacData+
ACER AUC ACC EER ACER AUC ACC EER ACER AUC ACC EER
CLIP-V 35.83 67.91 53.44 35.83 41.41 60.9 49.69 41.41 39.81 63.4 54.1 39.81
CLIP 39.32 63.95 55.65 39.32 40.58 62.75 50.38 40.58 39.36 64.08 51.92 39.36
CoOp(unified) 39.59 63.56 58.29 39.6 26.93 78.77 70.44 26.93 29.77 76.77 67.4 29.77
CoOp(specific) 39.39 63.89 58.71 39.39 28.62 77.5 69.9 28.62 30.3 75.9 68.49 30.3
HiPTune 32.88 69.34 60.51 32.88 20.16 88.53 86.82 20.23 25.93 83.05 73.82 25.94
UniAttackData-JFSFDB UniAttackData+-JFSFDB UniAttackData+UniAttacData+-JFSFDB
ACER AUC ACC EER ACER AUC ACC EER ACER AUC ACC EER
CLIP-V 36.16 64.97 70.81 36.16 51.65 48.75 47.88 51.65 45.12 56.26 50.52 45.12
CLIP 36.98 64.05 76.04 36.98 50.72 48.25 44.09 50.72 44.42 58.1 54.11 44.42
CoOp(unified) 40.02 64.12 58.53 40.02 43.55 59.51 54.28 43.55 38.21 65.84 59.44 38.21
CoOp(specific) 39.44 64.83 59.02 39.44 42.95 59.99 55.39 42.95 38.13 66.04 60.67 38.13
HiPTune 34.39 71.23 80.31 34.42 28.14 79.22 70.46 28.78 32.93 70.89 72.23 32.92
JFSFDB-UniAttackData UniAttackData+-UniAttackData JFSFDB+UniAttackData+-UniAttacData
ACER AUC ACC EER ACER AUC ACC EER ACER AUC ACC EER
CLIP-V 29.94 75.58 62.11 29.94 22.26 84.22 56.92 22.26 13.15 91.86 77.74 13.15
CLIP 29.27 80.12 67.63 29.27 22.38 84.11 75.76 22.38 16.59 90.14 82.5 16.59
CoOp(unified) 37.17 67.2 58.91 37.17 25.4 80.5 71.77 25.4 30.28 76.11 65.41 30.28
CoOp(specific) 41.4 61.95 57.39 41.4 26.37 79.32 70.68 26.37 27.34 79.19 67.3 27.34
HiPTune 21.36 85.59 74.39 31.56 18.62 90.75 88.64 18.63 8.55 94.67 90.26 8.56
TABLE VIII: Results on existing UAD datasets.
In the UAD-Benchmark, training and testing models across different datasets allow us to evaluate not only their generalization to various attack types but also their domain generalization capabilities. By examining the table, we can draw two key insights:
(1) As shown in the second column of the table, models trained on UniAttackData+ demonstrate strong anti-forgery capabilities that can be effectively transferred to other datasets. This highlights the robustness and generalizability of the knowledge learned from UniAttackData+.
(2) Baseline models show limited improvement, suggesting that we should design models specifically tailored to the characteristics of the dataset. Such models should be capable of fully leveraging the knowledge contained within the dataset, maximizing its utility for model training and evaluation.
V-D Results on JFSFDB
Face Spoofing Detection Face Forgery Detection
Intra-testing Cross-testing Intra-testing Cross-testing
Metrics Sep./ Joint Method SiW 3DMAD HKBU MSU 3DMask ROSE FF++ DFDC Celeb-DFv2 Average
MesoNet 95.73 100 68.87 85.33 44.54 68.05 65.54 54.72 65.1 71.99
Xception 99.88 100 94.47 90.62 63.88 90.15 98.21 64.86 84.27 87.37
MultiAtten 99.56 100 68.68 89.21 36.88 72.66 70.24 63.37 64.34 73.88
CDCN++ 100 99.56 94.51 82.65 47.66 76.76 97.53 56.65 78.34 81.51
DeepPixel 100 100 89.4 95.17 56.87 80.6 99.26 67.54 85.51 86.04
ResNet50 99.99 100 93.71 94.33 58.95 88.79 97.04 62.35 82.51 86.41
ViT 100 100 95.45 95.13 35.6 89.12 98.65 73.79 86.61 86.04
CLIP-V 99.99 100 96.85 97.29 76.62 93.6 93.05 82.09 75.4 90.54
CLIP 99.99 100 95.58 95.82 80.4 92.89 95.92 86.08 79.4 91.79
CoOp(Unified) 99.86 100 94.42 83.73 91.94 94.63 79.49 64.46 62.05 85.62
CoOp(Specific) 99.83 100 93.72 84.14 92.61 92.89 81.41 66.63 61.88 85.9
Sep. HipTune 99.97 100 98.62 98.62 95.87 98/62 100 89.72 89.63 93.12
MesoNet 92.55 97.94 78.42 84.01 45.87 70 56.66 50.57 55.03 72.56
Xception 99.8 100 97.62 91.56 75.49 79.38 96.69 63.69 78.87 87.01
MultiAtten 99.38 100 78.34 92.89 58.31 77.51 69 58.11 60 77.06
CDCN++ 99.99 98.89 96.8 78.93 61.54 76.39 95.79 56.42 76.92 82.41
DeepPixel 100 100 97.31 93.97 66.65 74.58 99.22 66.36 88.45 87.39
ResNet50 99.93 100 98.1 93.5 79.64 86.03 98.83 67.18 85.55 89.86
ViT 100 100 100 95.85 62.74 91.88 98.83 75.11 88.36 90.31
JFSFDB-Uni 100 100 99.96 97.99 67.1 95.57 99.12 75.9 89.76 91.71
CLIP-V 100 100 95.91 97.23 71.26 83.96 95.69 82.84 79.64 89.44
CLIP 100 100 97.48 96.39 86.13 81.07 94.83 84.06 78.99 90.83
CoOp(Unified) 99.71 100 95.82 81.58 89.27 93.62 78.55 68.56 64.88 85.78
CoOp(Specific) 99.78 100 95.33 80.68 92.82 93.34 76.42 67.58 65.06 85.67
AUC(%) Joint HipTune 100 100 100 99.62 95.43 97.82 100 90.88 93.35 97.46
TABLE IX: Results on JFSFDB AUC.
Face Spoofing Detection Face Forgery Detection
Intra-testing Cross-testing Intra-testing Cross-testing
Metrics Sep./ Joint Method SiW 3DMAD HKBU MSU 3DMask ROSE FF++ DFDC Celeb-DFv2 Average
MesoNet 11.67 0 35.12 21.9 55.82 36.81 38.93 46.16 38.78 31.69
Xception 1.14 0 11.31 18.57 42.3 18.13 5.54 38.92 24.07 17.78
MultiAtten 2.78 0 35.71 18.57 61.22 33.82 35.54 39.73 40.46 29.76
CDCN++ 0.14 3.33 11.9 22.8 52.76 32.01 7.5 44.75 28.19 22.6
DeepPixel 0 0 17.86 15.71 46.42 28.23 2.32 36.82 22.81 18.91
ResNet50 0.36 0 13.69 12.38 45.24 19.94 8.18 40.61 26.01 18.49
ViT 0.14 0 11.31 11.9 60.4 19.47 4.64 32.75 22.76 18.15
CLIP-V 0.49 0 7.74 8.58 29.56 13.17 13.66 25.43 29.89 14.28
CLIP 0.34 0 11.51 12.86 26.74 14.61 11.47 23.14 27.16 14.2
CoOp(Unified) 1.51 0 12.7 25.95 17.02 12.06 27.14 39.42 41.21 19.67
CoOp(Specific) 1.66 0 15.08 25.48 15.68 15.25 25.98 37.69 41.53 19.82
Sep. HipTune 0.35 0 4.58 6.56 11.57 12.82 1.89 20.86 28.77 9.71
Mesonet 14.8 6.67 28.57 22.86 54.76 36.18 44.29 49.81 46.94 33.88
Xception 1.71 0 9.52 17.14 22.21 28.82 9.29 40.75 29.16 17.62
MultiAtten 4.13 0 28.57 13.81 44.42 30.32 35.96 43.93 43.57 27.19
CDCN++ 0.28 6.67 10.71 31.9 43.83 33.19 8.39 44.58 29.68 23.25
DeepPixel 0.14 0 10.12 13.33 40.19 32.8 2.14 37.38 19.48 17.29
ResNet50 1.21 0 6.55 12.86 28.79 21.55 4.29 36.82 23.06 15.01
ViT 0 0 0 9.52 46.3 16.2 2.68 30.99 20.01 13.97
JFSFDB-Uni 0 0 0.6 6.67 43.6 11.64 2.32 30.6 18.8 12.69
CLIP-V 0 0 10.71 8.33 37.17 23.46 11.43 25.72 27.81 16.8
CLIP 0 0 8.13 10.24 23.58 26.86 12.14 23.76 27.34 15.4
CoOp(Unified) 0.02 0 11.11 28.57 17.77 14.39 28.57 36.69 39.84 19.66
CoOp(Specific) 1.54 0 11.11 29.76 13.82 14.48 31.43 37.22 40.31 19.96
EER(%) Joint HipTune 0 0 0 3.84 11.72 8.99 1.23 24.92 12.21 6.99
TABLE X: Results on JFSFDB-EER.
JFSFDB primarily demonstrated that training with a combination of physical and digital attack samples can significantly enhance model performance. This conclusion is also applicable to our HiPTune model. When trained and tested separately within the dataset, HiPTune achieved an EER of 9.71% and an AUC of 93.12%. However, after joint training, HiPTune gained access to a more diverse set of classification criteria, which it could adaptively select and integrate, resulting in a substantial performance improvement: a 4.34% increase in AUC and a 2.72% reduction in EER. Moreover, the model’s generalization capability was significantly enhanced.
V-E Results on UniAttackData
Protocol Methods ACER ACC AUC EER Protocol Methods ACER ACC AUC EER
ResNet-50 1.35 98.83 99.79 1.18 ResNet-50 34.6 53.69 87.89 19.48
ViT-B/16 5.92 92.29 97 9.14 ViT-B/16 33.69 52.43 83.77 25.94
Auxiliary 1.13 98.68 99.82 1.23 Auxiliary 42.98 37.71 76.27 32.66
CDCN 1.4 98.57 99.52 1.42 CDCN 34.33 53.1 77.46 29.17
FFD 2.01 97.97 99.57 2.01 FFD 44.2 40.43 80.97 26.18
CLIP-V 0.46 99.55 99.95 0.46 CLIP-V 11.38 85.93 95.73 11.38
CLIP 0.58 97.4 99.93 0.58 CLIP 15.27 79.83 92.46 15.27
CoOp(unified) 7.94 90.96 97.9 7.94 CoOp(unified) 33.29 63.89 71.58 33.29
CoOp(specific) 8.49 90.98 97.45 8.49 CoOp(specific) 35.67 60.9 68.41 35.67
UniAttackDetection 0.52 99.45 99.96 0.53 UniAttackDetection 22.42 67.35 91.97 15.72
1 HiPTune 0.33 99.98 99.99 0.36 2 HiPTune 10.38 88.62 97.82 10.39
TABLE XI: Results on UniAttackData.
In the UniAttackData dataset, where the data distribution is relatively simple, our HiPTune model achieved outstanding performance. Leveraging the guidance function of prompts, HiPTune also delivered excellent results in cross-attack protocols, demonstrating its robustness and adaptability across different attack types.
V-F Abalation Study
To verify the effectiveness of the theory proposed by HiPTune and to analyze the influence of various settings, we conduct several ablation studies on protocol $P_{2}$ of HiFAB. These studies examine the depth of the tree, length of prompts, different selection strategies, various gate functions, and the contribution of each component. Additional experiments and detailed analyses of each element are provided in the Supplementary Materials.
Depth of Tree. The depth of the VP-Tree plays a crucial role in helping models understand fine-grained classification criteria, which enhances their ability to handle unseen advanced attacks during the integration process. Models such as CoOp, which lack a hierarchical structure, exhibit poor performance as they rely on a single, rigid classification criterion. In contrast, increasing the depth of the VP-Tree allows the model to ”draw” the line between live and fake faces adaptively and accurately, demonstrating the importance of a hierarchical approach.
Length of Prompts. The length of prompts determines the scope of learnable parameters, impacting both accuracy and efficiency. To explore this balance, we conduct several experiments to assess the effect of different prompt lengths. Using more parameters can capture more detailed classification-related information, enhancing model performance.
Component ACER(%) AUC(%) EER(%) ACC(%)
level-0 (CoOp) 14.20 92.57 14.26 85.44
level-1 9.78 94,26 10.12 90.12
level-2 5.78 96.88 5.45 95.08
level-3 3.26 98.78 3.67 95.79
TABLE XII: The influence of depth of the VP-Tree.
Component ACER(%) AUC(%) EER(%) ACC(%)
16 4.24 95.79 4.26 92.39
20 4.08 96.03 4.03 93.40
29 3.45 96.34 3.44 94.67
40 3.48 97.62 3.90 94.23
60 3.26 98.78 3.67 95.79
TABLE XIII: The impact on the length of prompts.
VI Conclusion
In this paper, we first present a large-scale Unified Face Attack Benchmark (HiFAB), which is the most extensive and comprehensive UAD dataset to date. It includes $14$ kinds of physical attacks, $40$ digital attacks, and $697,347$ videos in total. Every attack is applied to $2,875$ identities to reduce the ID noise. Furthermore, we introduce a novel Hierarchical Prompt Tuning framework (HiPTune), that adaptively integrates classification criteria from various semantic spaces, allowing tailored classification criteria for each sample. Finally, we conduct several experiments that evaluate the model’s performance in real-world scenarios.
VII Disscussion and Future Work
In this work, we explore the application of prompt tuning in the visual domain while acknowledging the value of examining its adaptability to its original text-based purpose. The hierarchical and semantic capabilities of text-based prompt tuning could provide valuable insights for visual applications, warranting further investigation. Furthermore, the suboptimal performance of baseline models under protocol $P_{1}$ may be partially attributed to the absence of a fine-tuning process on the testing datasets. This lack of adaptation limits the models’ ability to generalize to unseen attack types and domain variations. Further discussion on the impact of fine-tuning in bridging this gap will be presented.

### v3

@@
 Extensive experiments on 12 datasets have verified the potential to inspire further innovations in the UAD field.
+The UniAttackDataPlus dataset can be found at https://sites.google.com/view/face-anti-spoofing-challenge/dataset-download/uniattackdata-iccv2025, and the codes can be found at https://github.com/HaochengYUAN-Allen/UniAttackDataPlus-HiPTune.
 Index Terms:
-Unified Face Attack Detection, Face Anti-Spoofing, Face Forgery Detection, Prompt Tuning.
-I Introduction
+Unified Face Attack Detection, Face Anti-Spoofing, Face Forgery Detection, Benchmark, Prompt Tuning.
+1 Introduction
 Face data are widely used in fields such as facial payment, mobile unlocking, and access authentication, due to the discriminative and easily obtainable facial biometric features.
@@
 As shown in the upper left of Fig.
-2 (a), attackers can deceive face recognition systems by using physical media to impersonate others before information collection—a method known as physical attacks or presentation attacks.
+1 (a), attackers can deceive face recognition systems by using physical media to impersonate others before information collection, a method known as physical attacks or presentation attacks.
 Additionally, attackers may employ digital forgery (commonly referred to as deepfake) to edit or manipulate collected images for fraudulent purposes.
-To address the aforementioned threats, Face Anti-Spoofing (FAS), also known as Presentation Attack Detection (PAD), and Face Forgery Detection (FFD) have been proposed to ensure the security of facial data.
+To address the aforementioned threats, Face Anti-Spoofing, also known as Presentation Attack Detection (PAD), and Face Forgery Detection (FFD) have been proposed to ensure the security of facial data.
 Essentially, both methods are formulated as binary classification tasks, where the model determines the authenticity of an image by outputting a probability indicating whether it is real or fake.
-Researchers have proposed a wide variety of attack methods and have constructed multiple databases to advance the development of defense algorithms.
-Common physical presentation attacks are generally categorized into two types: 2D attacks (such as printed paper, fabric, screen replays) and 3D attacks (including 3D masks and silicone head models).
-Most models achieve authenticity classification by analyzing image texture details, focusing on inconsistent edges, screen moiré patterns, and other indicators.
-However, their generalization performance has been challenged.
-Consequently, domain generalization protocols such as ICMO and SCW have been proposed to evaluate model generalization.
-For instance, [1] utilized X, [2] leveraged X, and [3] employed X, while recent studies [4], [5], [6] have achieved notable results using vision-language models by employing X for X.
-Digital attack methods can be broadly classified into three categories: editing, adversarial attacks, and generative attacks.
-Digital editing attacks aim to modify image attributes (e.g., changing facial expressions, age) or perform face swapping (replacing the current face with another), video-driven manipulation (using a video to drive facial expressions), and more.
-Adversarial attacks can be divided into pixel-level and semantic-level attacks, which either modify or erase relevant pixels or alter high-dimensional semantics to manipulate the original facial data.
-Digital generative attacks, the most recent and innovative type, involve generating entirely new facial data using generative models.
-To combat digital attacks, researchers typically focus on detecting inconsistencies in pixel-level and semantic-level information to achieve robust classification performance.
-Figure 1: The dashed lines represent the classification criteria that separate real and fake faces in the feature space.
-(a) Existing models are trained to draw a single ”line” in the feature space, leading to unpromising classification results.
-(b) Our HiPTune can adaptively integrate the classification criteria from different semantic spaces to achieve robust detection towards various attacks.
-Different colors represent distinct classification criteria.
 Figure 1: (a) The Physical and Digital Attack construction approaches.
@@
 (b) Our UniAttackData+ is constructed in an explainable hierarchical way.
-The coarse-to-fine categories provide researchers with prior knowledge which is helpful for the model designs.
-(b) A comparison of HiFAB with existing FAS, FFD, and UAD datasets shows that HiFAB surpasses all others in terms of attack types and video quantity.
+The coarse-to-fine categories provide researchers with prior knowledge, which is helpful for the model designs.
+(c) A comparison of HiFAB with existing FAS, FFD, and UAD datasets shows that HiFAB surpasses all others in terms of attack types and video quantity.
 The radius of the circle represents the scale of the dataset, which is the number of videos contained in the dataset, corresponding to the horizontal axis of the figure.
@@
 The number following each category name indicates the number of attack types included.
-Existing PAD and FFD models are typically trained and deployed separately, which inevitably reduces their robustness against unknown attacks and increases deployment complexity.
-When confronted with unknown attack samples or hybrid attack types, models trained independently exhibit poor robustness against such attacks.
+Researchers have proposed a wide variety of attack methods and have constructed multiple databases to advance the development of defense algorithms.
+Common physical presentation attacks are generally categorized into two types: 2D attacks (i.e., print attack [1] and video attack [2]), and 3D attacks (i.e., mask [3] and fake head [4]).
+Most models [5, 6, 7, 8, 9, 10, 11] achieve authenticity classification by analyzing image texture details, focusing on inconsistent edges, color distortion, screen moiré patterns, and mask pouring traces.
+However, their generalization severely deteriorates when faced with unknown domain datasets.
+Currently, domain generalization based PAD algorithms  have become a hot research topic, such as UDG-FAS [12], SA-FAS [13], IADG [14], FLIP-MCL [15], GAC-FAS [16], and TTDG-V [17].
+Digital attack methods can be broadly classified into three categories: editing, adversarial attacks, and generative attacks.
+Digital editing attacks aim to modify image attributes [18, 19] (e.g., changing facial expressions, age) or perform face swapping [20, 21] (e.g., replacing the current face with another), video-driven manipulation [22, 23] (e.g., using a video to drive facial expressions).
+Adversarial attacks can be divided into pixel-level [24, 25, 26, 27] and semantic-level attacks [28, 29, 30], which either modify or erase relevant pixels or alter high-dimensional semantics to manipulate the original facial data.
+Digital generative attacks [31, 32, 33], the most recent and innovative type, involve generating entirely new facial data using generative models.
+To combat digital attacks, researchers [34, 35, 36] typically focus on detecting inconsistencies in pixel-level and semantic-level information to achieve robust classification performance.
+Figure 2: The dashed lines represent the classification criteria that separate real and fake faces in the feature space.
+(a) Existing models are trained to draw a single ”line” in the feature space, leading to unpromising classification results.
+(b) Our HiPTune can adaptively integrate the classification criteria from different semantic spaces to achieve robust detection towards various attacks.
+Different colors represent distinct classification criteria.
+Existing PAD and FFD models are usually trained and deployed independently, leading to weak robustness against unknown or hybrid attacks and increased deployment complexity.
 A unified model that can leverage the advantages of different attack types while maintaining a lightweight parameter size is urgently needed.
-We attribute the lack of such a unified model to two primary reasons:
+We identify two main reasons for the absence of such a solution:
 (1) The lack of a comprehensive dataset that provides the model with a broader learning space.
-Although there are currently three unified face attack detection datasets (JFSFDB, Grandfake, UniAttack), they simply mix existing datasets that are tailored to specific attack types.
+Although there are currently three unified face attack detection datasets (i.e., GrandFake [37] and JFSFDB [38]), they simply mix existing datasets that are tailored to specific attack types.
 This approach introduces substantial noise, such as variations in style, background, and identity, making it difficult to accurately evaluate the model’s superiority and generalization capability.
@@
 1 (b), their sizes are relatively small, containing limited attack types, and many of them are not publicly accessible, making them difficult to utilize effectively.
-To mitigate these issues, we previously proposed UniAttackData, a dataset where a total of 14 attacks (2 physical and 12 digital) are applied to the same face, creating an identity-consistent face attack dataset.
-However, its size and the diversity of attack types remain insufficient, resulting in models that tend to overfit the training set and perform poorly when encountering previously unseen attack types.
+To address these challenges, we previously introduced UniAttackData [39], a unified dataset featuring 14 attack types, including 2 physical and 12 digital attacks, all applied to the same identity.
+This design ensures identity consistency across diverse attacks, providing a valuable benchmark for training and evaluating generalized UAD models.
+However, its size and the diversity of attack types remain insufficient, especially lacking the most advanced generative face attack methods [31, 32, 33] currently available.
 (2) The absence of a reliable classification criterion that can accurately categorize each sample.
 Existing methods typically aim to identify a single classification criterion within a single semantic space, as illustrated in Fig.
-1 (a), where the model is trained to discover a single ’line’ that separates real and fake faces.
+2 (a), where the model is trained to discover a single ‘line’ that separates real and fake faces.
 However, in the context of UAD, this single classification criterion cannot robustly classify all sample types.
@@
 The feature extractor struggles to effectively cluster them within a common space, making it difficult for a single classification criterion to determine the class of each sample accurately.
-To address the dataset limitations mentioned above, we adopted the identity-consistent approach of UniAttackData and developed UniAttackDataPlus (UniAttackData+), which is currently the largest publicly available unified face attack detection dataset.
+To address the dataset limitations mentioned above, we adopted the identity-consistent approach of UniAttackData [39] and developed UniAttackDataPlus (UniAttackData+), which is currently the largest publicly available unified face attack detection dataset.
 It features the most extensive variety of attack types, minimal interference, and highly structured characteristics, making it a robust and comprehensive resource for training and evaluating face attack detection models.
-We collected data from $2,875$ participants representing three distinct ethnic groups (African, East Asian, and Central Asian), capturing $18,250$ authentic videos under various lighting conditions, backgrounds, and acquisition devices.
+We collected data from $2,875$ participants representing three distinct ethnic groups (i.e., African, East Asian, and Central Asian), capturing $18,250$ authentic videos under various lighting conditions, backgrounds, and acquisition devices.
 For each participant, we applied $54$ different attack methods, including $14$ physical attacks and $40$ digital attacks, resulting in a total of $679,097$ forged videos, which is $32$ times more than the existing dataset.
-From each video, we randomly extracted at least $30$ frames to construct the dataset; samples can be found in Fig.
+The statistical comparison is shown in Fig.
 1 (c).
+From each video, we randomly extracted at least $30$ frames to construct the dataset.
 The diverse acquisition backgrounds of authentic faces provide a more varied dataset, encouraging models to focus on discriminative features rather than irrelevant factors.
@@
 Most importantly, UniAttackData+ applies a full range of attack methods to each unique identity, covering almost all known attack types.
-This not only prevents models from achieving high performance through ’identity memorization’ but also enhances the dataset’s applicability in real-world deployment scenarios.
-Additionally, our dataset is structured in a hierarchical manner, with all attack types categorized from coarse to fine, as depicted on the right side of Fig.
-2 (a).
+This not only prevents models from achieving high performance through ‘identity memorization’, but also enhances the dataset’s applicability in real-world deployment scenarios.
+Additionally, our dataset is structured in a hierarchical manner, with all attack types categorized from coarse to fine, as depicted in Fig.
+1 (d).
 This hierarchical structure provides researchers with a more interpretable framework for designing corresponding algorithms.
-For instance, our baseline method, HiPTune, leverages a similar hierarchical approach to develop a prompt-tuning method.
-Furthermore, we have designed four protocols to evaluate model performance under various scenarios, including unknown identities and unseen attack methods.
-We also introduced a cross-dataset testing benchmark to assess model generalization.
-Compared to previous datasets that simply merged existing datasets, UniAttackData+ offers a robust benchmark for the face data security community in a more interpretable hierarchical format, significantly promoting the development of more advanced algorithmic models.
-To address the issue of insufficient robustness caused by a single classification criterion in unified attack detection, we propose a hierarchical prompt strategy named Hierarchical Prompt Tuning (HiPTune).
+To address the issue of insufficient robustness caused by a single classification criterion in UAD, we propose a hierarchical prompt strategy named Hierarchical Prompt Tuning (HiPTune).
 HiPTune is designed to explore multiple classification criteria across different semantic spaces and adaptively integrate the most appropriate criterion for each sample.
-Specifically, following the hierarchical structure of our dataset (as shown on the right side of Fig.
-2 (a)), we construct a Visual Prompt Tree (VP-Tree) as a carrier of multiple classification criteria.
-These learnable parameters, referred to as visual prompts, are concatenated with image tokens and fed into a fixed image encoder for training.
-These prompts serve as guiding symbols that direct the encoder to extract critical discriminative features, which we define as our ’classification criteria.’ The visual prompts are organized in a tree structure, progressing from coarse to fine, where each level represents a different granularity of classification criteria, and each leaf node corresponds to a specific criterion that is robust enough to distinguish certain fake samples from real faces.
-By inserting these visual prompts into different depths of the Transformer blocks in a top-down manner, we ensure that the model first learns to capture macro-level differences, achieving comprehensive detection of forgeries.
-As the feature extraction process deepens, finer-grained prompts focus on local details, enabling precise forgery discrimination.
+Specifically, following the hierarchical structure of our dataset (as shown the Fig.
+1 (d)), we first construct a Visual Prompt Tree (VP-Tree) as a carrier of multiple classification criteria.
+The visual prompts are organized in a tree structure, progressing from coarse to fine, where each level represents a different granularity of classification criteria, and each leaf node corresponds to a specific criterion that is robust enough to distinguish certain fake samples from real faces.
 Shown as Fig.
@@
 This cross-modal integration strategy broadens the semantic coverage of classification criteria, ensuring classification accuracy from multiple dimensions.
-Subsequently, we calculate the self-attention scores of the input samples, guiding the model to adaptively select the appropriate classification criteria at different levels for each sample.
-This adaptive strategy preserves model autonomy, avoiding excessive dependence on manual knowledge while uncovering potential connections between different attacks.
-It enables more robust detection tailored to each input sample.
+Finally, we ultilize a Adaptive Prompt Pruning (APP) mechanism that calculate the cross-attention scores between the input samples and the prompts, guiding the model to adaptively select the appropriate classification criteria at different levels for each sample.
 To summarize, our contributions are as follows:
-  • We create UniAttackData+, the largest publicly available unified face attack detection dataset, encompassing $54$ attacks per identity in an explainable hierarchical manner.
-  • We proposed the Hierarchical Prompt Tuning (HiPTune) framework, which adaptively integrates classification criteria from different semantic spaces to establish a more robust criterion for each input sample.
-  • We propose four intra-dataset protocols and nine cross-dataset benchmarks for the Unified Face Attack Detection community that promote the development of the detection model.
-II Related Works
-II-A Unified Face Attack Detection
-II-A1 Datasets
-Recently, several datasets, including Grandfake [1], JFSFDB [2], and UniAttack [3], have been introduced to address the challenges of UAD.
-These datasets combine existing PAD and FFD datasets to encompass a variety of attack types within a single dataset.
-Additionally, UniAttackData [4] was proposed as a UAD dataset with identity (ID) consistency, where each ID includes a complete range of attack types, thereby reducing ID and domain noise.
-The current UAD dataset not only lacks emerging generative forgery face samples but also outdated forgery algorithms corresponding to digitally edited faces, making it impossible to train a high-quality joint attack detection model.
-Our UniAttackData+ addresses this gap by incorporating a broader range of modern attacks and structuring them in an explainable manner, rather than merely aggregating datasets.
-II-A2 Approaches
-For the UAD task, Deb et al.
-[1] introduced a multi-task learning approach that treats each attack type as a separate detection task based on feature similarity.
-Yu et al.
-[2] focused on improving generalization by utilizing a multi-modal approach for joint task training.
-Fang et al.
-[4] leveraged the Vision-Language Model to capture both unified and specific knowledge, while Cao et al.
-[3] aimed to enhance face anti-spoofing by promoting the learning of compact, fundamental features of genuine faces.
-However, these methods rely on a single, generalized classification criterion, which often proves inadequate when faced with advanced attacks.
-In contrast, our HiPTune framework adaptively integrates the most suitable prompts from various semantic spaces, enabling more effective classification.
-II-B Prompt Learning
-PREFER [5] enhances prompt ensemble learning with a feedback mechanism for iterative refinement and prompt bagging, improving stability and performance across tasks.
-VPT [6] extends prompt learning from NLP to visual tasks.
-S-Prompt [7] clusters domain samples using K-NN, while C-Prompts [8] selects prompts from different domains for robust Domain Incremental Learning.
-CoOp [9] demonstrated that replacing manual prompts in CLIP with learnable contexts enhances performance via few-shot learning.
-CoCoOp [10] further adapts prompts to new domains with a Meta-Net, enhancing generalization.
-However, existing methods overlook hierarchical prompt relationships, which capture both global and detailed knowledge.
-HiPro [11] organizes tasks hierarchically but lacks clarity between tasks.
-Our proposed HiPTune is able to adaptively select and integrate classification criteria from different semantic spaces by the VP-Tree, ensuring that the model not only learns to capture macro-level differences but also focuses on local details for the precise forgery discrimination.
+  • We built on the identity-consistent design of UniAttackData [39] to create UniAttackData+, the largest UAD dataset to date, featuring $54$ attack types and over $679,000$ videos from $2,875$ participants.
+Collected from $2,875$ participants across three ethnic groups, the dataset includes 18,250 real videos under diverse conditions, with at least 30 frames sampled per video for training and evaluation.
+  • We propose HiPTune, a hierarchical prompt tuning strategy that leverages a VP-Tree and a APP to organize multiple classification criteria from coarse to fine, enabling adaptive selection for each sample.
+By integrating a DPI module with self-attention, HiPTune enhances semantic richness and cross-modal alignment, allowing the model to balance global generalization and local discrimination for improved face attack detection.
+  • We propose four intra-dataset protocols and nine cross-dataset benchmarks to promote the development of UAD community.
+TABLE I: Comparison of existing PAD and FFD datasets.
+The scale of our UniAttackData+ is far more than the types in the existing PAD and FFD datasets.
+‘(I)’ indicates that the dataset only contains images.
+Dataset # Attack Types # Videos
+CASIA-FASD [40] phy.2 600
+Replay-Attack [2] phy.3 1,200
+MSU-MFSD [41] phy.3 440
+HKBU-MARs V2 [42] phy.2 1,008
+OULU-NPU [43] phy.4 5,940
+SiW [5] phy.6 4,620
+Rose-Youtu [44] phy.4 4,000
+CelebA-Spoof [45] phy.4 625,537(I)
+3DMask [46] phy.3 1,152
+SiW-Mv2 [47] phy.14 1,700
+HiFiMask [3] phy.3 54,600
+WMCA [4] phy.7 6,716
+CASIA-SURF [1] phy.6 21,000
+PADISI-Face [48] phy.9 121,740(I)
+DFD [49] digi.5 3,431
+DFDC [50] digi.8 128,064
+FaceForensics++ [51] digi.4 5,000
+Celeb-DF [52] digi.1 6,229
+DeeperForensics-1.0 [53] digi.1 60,000
+FFIW [54] digi.3 20,000
+ForgeryNet [55] digi.15 221,247
+UniAttackData+ phy.14, digi.40 697,347
+2 Related Works
+2.1 Face Attack Datasets
+2.1.1 PAD Datasets
+CASIA-FASD [40] comprises 50 subjects, with each subject having 12 videos (3 genuine and 9 fake).
+The videos are captured in three different imaging qualities: low, normal, and high.
+Two types of face attacks are included in the dataset: photo and video attacks.
+Replay-Attack [2], comprises short video recordings of both real-access and attack attempts on 50 clients in 1,200 videos.
+It includes three types of attacks: two kinds of replayed faces and one kind of printed face.
+MSU-MFSD [41] comprises 440 videos across 35 subjects, employing two mobile devices to investigate three types of spoof attacks: 1 printed photo, and 2 video replay.
+HKBU-MARs V2 [42] contains 1,008 videos from 12 subjects and 2 types of 3D masks: Thatsmyface and REAL-F masks.
+OULU-NPU [43] contains 5,940 high-resolution videos from 55 subjects, including 1,980 real-access and 3,960 attack videos.
+Captured with six mobile front cameras over three sessions, it features varying lighting and backgrounds, and includes 2 printers and 2 displays.
+SiW [5] contains 4,620 30-fps videos from 165 subjects, where each subject has 6 spoof types: 2 print, and 4 replay.
+Live videos are recorded across four sessions featuring variations in head movement, facial expression, and illumination.
+Rose-Youtu [44] contains over 4,000 videos of 25 subjects captured under varied illumination, camera models, and three spoofing attacks (printed paper, video replay, and masking).
+CelebA-Spoof [45] is a richly annotated dataset with 625,537 images from 10,177 subjects.
+It includes four types of spoofing attacks, i.e., print, paper-cut, replay, and 3D mask, captured under four illumination conditions and two environmental settings, i.e., indoor and outdoor.
+SiW-Mv2 [47] has 785 videos from 493 live subjects, and 915 spoof videos from 600 subjects.
+Among these spoof videos, it has 14 spoof attack types, spanning from typical 2D spoof attacks (e.g., print and replay), various masks, different makeups, and physical material coverings.
+3DMask [46] contains 288 live videos and 864 3D mask videos from 48 subjects.
+Uniquely, it includes outdoor scenes with challenging lighting and considers three mask decoration types.
+HiFiMask [3] is a realistic masks dataset, with a total of 54,600 videos recorded from 75 subjects with 3 kinds of high-fidelity masks.
+WMCA [4] includes a diverse range of 2D and 3D presentation attacks that can be grouped into 7 categories.
+It provides 6,716 videos with multi-channel data, i.e., color, depth, near-infrared, and thermal.
+CASIA-SURF [1] is the first large-scale multimodal PAD dataset, which consists of 1,000 subjects with 21,000 videos, and each sample has 3 modalities (i.e., RGB, Depth, and IR) and 6 print attacks.
+PADISI-Face [48] is a realistic and challenging dataset, and includes 121,740 frames from 9 attack types.
+2.1.2 FFD Datasets
+DFD [49] dataset, developed by Google in partnership with Jigsaw, contains 3,431 high-quality videos of actors speaking in 5 deepfake approaches.
+DFDC [50] is the largest publicly available face swap video dataset, with 128,064 video clips, produced with 8 Deepfake, GAN-based, and non-learned methods.
+FaceForensics++ [51] contains 4,000 videos, and manipulations created with 4 methods, namely, Face2Face, FaceSwap, DeepFakes, and NeuralTextures.
+Celeb-DF [52] contains 590 real videos, and 5,639 high-quality DeepFake videos of celebrities generated with a synthesis algorithm.
+DeeperForensics-1.0 [53] contains 60,000 videos constituted by a total of 17.6 million frames, where the 10,000 fake videos are generated by a proposed face swapping framework.
+FFIW [54] comprises 10,000 real and high-quality forgery videos, with an average of three human faces in each frame, and only a small subset has been manipulated with one of three face swapping methods.
+ForgeryNet [55] is the largest publicly available deepfake dataset, containing 2.9 million images and 221,247 videos with unified annotations across image- and video-level tasks.
+It supports four major tasks, i.e., image and video forgery classification, spatial forgery localization, and temporal forgery localization, spanning 15 manipulation methods and 36 perturbations.
+2.1.3 UAD Datasets
+GrandFake [37] is the first UAD dataset developed to resolve the dilemma wherein a single model struggles to handle all three categories of attacks concurrently.
+It consists of 25 face attacks from 3 attack categories, i.e., Adversarial Faces, Digital Manipulation, Physical Spoofs.
+JFSFDB [38] dataset consists of 9 existing datasets, including 6 PAD datasets and 3 FFD datasets.
+However, the limited diversity of attack types and data volume in the GrandFake and JFSFDB datasets, coupled with the absence of identities shared across all attack types, constrains the model’s generalization and makes it prone to overfitting identity-specific features.
+Based on the CASIA-SURF CeFA dataset [56], UniAttackData [39] introduces 12 types of digital forgery on real faces, including 6 adversarial attacks and 6 attribute editing attacks.
+Although UniAttackData ensures that each identity includes all types of attacks, it is limited in the diversity of physical and digital forgeries, lacking high-fidelity mask attacks and advanced generative techniques.
+2.2 Face Attack Detection Methods
+UniFAD [37] is the first UAD method that combines k-means clustering and multi-task learning to jointly learn representations for coherent attack types while handling uncorrelated ones separately.
+UniAttackDetection [39] is a vision-language-based UAD method that integrates teacher-student prompts, unified knowledge mining, and sample-level prompt interaction to effectively capture both general and sample-specific attack semantics.
+La-SoftMoE [57] leverages a flexible Mixture of Experts with self-adaptive weighting to effectively handle sparse and diverse attack data across complex feature spaces.
+MoAE-CR [58] combines Mixture-of-Attack-Experts for fine-grained feature extraction with class-aware regularization modules to enhance intra-class compactness and inter-class separability for improved detection of diverse attacks.
+FA3-CLIP [59] combines spatial and frequency features through dual-stream fusion and attack-agnostic prompt learning to generate generalized live/fake representations, enabling robust detection across diverse physical and digital attacks.
+TABLE II: Comparison of existing UAD datasets.
+UniAttackData+ surpasses all other datasets in scale and diversity.
+‘V’ and ‘I’ indicate video- and image-based counts, respectively.
+The numbers following each attack type in the ‘# Catagories’ row represent the number of algorithms included in that category.
 Dataset Attack Type (each ID) # Datasets / Data # ID Physical Attacks Digital Attacks
@@
 HiFiMask 40,950 (V) Generation (7) 113,712 (V)
-TABLE I: Comparison of existing UAD datasets.
-UniAttackData+ surpasses all other datasets in scale and diversity.
-”V” and ”I” indicate video- and image-based counts, respectively.
-The numbers following each attack type in the ”# Catagories” row represent the number of algorithms included in that category.
-Dataset # Attack Types # Videos
-CASIA-SURF phy.6 21,000
-CeFA phy.3 23,538
-HiFiMask phy.5 54,600
-SiW-v2 phy.14 1,700
-MSU-MFSD phy.2 440
-CASIA-FASD phy.3 600
-ReplayAttack phy.2 1200
-OULU-NPU phy.2 5940
-FaceForensics++ digi.5 6,000
-DeeperForensics-1.0 digi.1 60,000
-DFDC digi.8 128,064
-ForgeryNet digi.15 221,247
-UniAttackData+ phy.10, digi.40 697,347
-TABLE II: Comparison of existing PAD and FFD datasets. The scale of our UniAttackData+ is far more than the types in the existing PAD and FFD datasets. The ”#” represents ”the number of”.
-III UniAttackDataPlus
-III-A Advantages of UniAttackData+
-Our UniAttackData+ dataset addresses five critical issues in current UAD datasets, i.e., insufficient sample size, limited attack coverage, and lack of interpretability, providing a comprehensive and valuable benchmark for the UAD community.
+3 UniAttackDataPlus
+3.1 Advantages of UniAttackData+
+Our UniAttackData+ addresses three critical issues in current UAD datasets, i.e., insufficient sample size, limited attack coverage, and lack of interpretability, providing a comprehensive and valuable benchmark for the UAD community.
 Scale.
@@
 Detailed data is presented in Tab.
-I and Tab.
-II.
+II and Tab.
+I.
 Our UniAttackData+ shows its superiority in separate PAD and FFD datasets, and also in the existing UAD datasets.
-The large dataset ensures a comprehensive and complete learning space, guaranteeing the effectiveness of detection models from a data perspective.
+It ensures a comprehensive and complete learning space, guaranteeing the effectiveness of detection models from a data perspective.
 Attack Types.
@@
 All the attack methods included in our dataset were proposed after 2021, incorporating the latest digital generation attacks.
-UniAttackData+ ensures the data’s applicability in real-world scenarios, significantly expanding the range of attack types for model training and reducing vulnerability to emerging advanced attacks.
+UniAttackData+ promotes real-world generalizability by significantly expanding attack diversity, enabling more comprehensive training and reducing model susceptibility to novel advanced attacks.
 Interpretability.
@@
 This simplistic combination hinders the model from focusing on genuine forgery cues and greatly reduces model robustness in real-world applications.
-In contrast, UniAttackData+ preserves the identity consistency from UniAttackData while structuring all attack types hierarchically.
-Each attack method and sample is categorized from coarse to fine and labeled accordingly.
-This construction method makes full use of prior knowledge, offering researchers an actionable perspective to address UAD challenges.
-The hierarchical label collection and setup provide a more comprehensive experimental and analytical space, enabling deeper and more interpretable analyses of detection models and algorithms.
-III-B Collection of UniAttackData+
+Unlike previous datasets, UniAttackData+ maintains the identity consistency of UniAttackData [39] and further organizes all attack types into a hierarchical taxonomy, enabling each sample to be labeled from coarse to fine levels.
+The hierarchical labeling scheme offers a richer experimental and analytical framework, facilitating more interpretable and fine-grained evaluation of detection models.
+3.2 Collection of UniAttackData+
 Live Face source.
-The live faces in the UniAttackData+, enhancing diversity in terms of face identity, angle, expression, and ethnicity, are sourced from three public datasets: CASIA-SURF [12], which includes $1,000$ subjects; CASIA-SURF CeFA [13] (CeFA), comprising 1,800 subjects from three ethnicities - African, East Asian, and Central Asian, with $600$ subjects each ethnicity; and CASIA-SURF HiFiMask [14] (HiFiMask), which contains $75$ subjects.
+The live faces in the UniAttackData+, enhancing diversity in terms of face identity, angle, expression, and ethnicity, are sourced from three public datasets: CASIA-SURF [1], which includes $1,000$ subjects; CASIA-SURF CeFA [56] (CeFA), comprising 1,800 subjects from three ethnicities, i.e., African, East Asian, and Central Asian, with $600$ subjects each ethnicity; and CASIA-SURF HiFiMask [3] (HiFiMask), which contains $75$ subjects.
 These varied sources significantly enhance the dataset’s diversity across multiple dimensions.
 We merge three existing datasets for the live face data to obtain $2,875$ distinct identities.
-For CASIA-SURF, the Intel RealSense SR300 camera captures collectors turning left or right, moving up or down, and approaching or moving away from the camera.
-The same collection method is used for the CASIA-SURF CeFA dataset.
-HiFiMask contains multiple lighting conditions, such as white, green, and periodic three-colors.
-Additionally, supplementary light sources from six directions are used to simulate real-world application scenarios.
+CASIA-SURF and CeFA use the RealSense SR300 to capture head movements and distance variations.
+HiFiMask simulates real-world lighting with white, green, and tri-color cycles, plus six-directional supplementary lights.
+Preprocessing.
 For dataset preprocessing, we apply MTCNN [60] for face detection and cropping on the real face data from the original datasets, with the scale parameter set to 1.3.
 We then manually assign class labels, attack type labels, and hierarchical labels to facilitate research.
-III-C Spoofing Approach
+3.3 Spoofing Approach
 To ensure diversity in physical attack methods within the UniAttackData+ dataset, we incorporate $14$ physical face spoofing techniques, including $11$ 2D presentation attacks and $3$ 3D mask attacks.
-The 2D attacks consist of $6$ paper-cutout variations and $2$ print attack types, as well as $3$ video-replay attacks to enrich the dataset diversity further.
-For 3D mask attacks, UniAttackData+ includes high-fidelity masks made from three different materials: transparent, plaster, and resin.
+See from the Tab.
+3, The 2D attacks consist of $6$ paper-cutout variations and $2$ print attack types, as well as $3$ video-replay attacks.
+For 3D mask attacks, UniAttackData+ includes high-fidelity masks made from $3$ different materials: transparent, plaster, and resin.
 These masks are recorded under varied lighting conditions and with multiple imaging devices for added variation.
@@
 All masks are worn on a real person’s face and uniformly dressed to avoid the algorithm looking for clues outside the head area.
-Category Type Methods # Videos Total
-2D Print Cloth-Indoor 1,800 3,600 55,950
-Cloth-Outdoor 1,800
-Replay pic-phone 1,800 5,400
-screen-phone 1,800
-screen-rs 1,800
-Cutouts Cutout1 1,000 6,000
-Cutout2 1,000
-Cutout3 1,000
-Cutout4 1,000
-Cutout5 1,000
-Cutout6 1,000
-3D Transparent 13,650 40,950
-Plaster 13,650
-Resin 13,650
-TABLE III: The detailed statistics of physical attacks in UniAttackData+. The ”#” represents ”the number of”.
-III-D Forgery Approach
-III-D1 Digital Manipulation Attacks
-Digital Manipulation Attacks in UniAttackData+ encompasses $17$ specific methods grouped into three types of attacks: (1) Face Swapping [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], where identity features of a source face, $I_{\text{source}}$, are transferred into a target face, $I_{\text{target}}$, producing a composite image, $I_{\text{composite}}=f(I_{\text{source}},I_{\text{target}})$.
+Figure 3: This diagram categorizes attacks into four distinct types: Physical Attacks, Digital Manipulation, Digital Adversarial and Digital Generation.
+Physical Attacks comprise 14 subtypes, Digital Manipulation includes 17 subtypes, the Digital Adversarial group contains 16 subtypes, and the Digital Generation group has 7 subtypes.
+We can clearly see two advantages of datasets containing forgery methods: (2) rich and diverse types of attacks; (2) Determine the proportion based on the commonness and progressiveness of attack types.
+3.4 Forgery Approach
+3.4.1 Digital Manipulation Attacks
+As shown in Tab.
+3, Digital Manipulation Attacks in UniAttackData+ encompasses $17$ specific methods grouped into three types of attacks: (1) Face Swapping [20, 61, 21, 62, 63, 64, 65, 66, 67, 68, 69], where identity features of a source face, $I_{\text{source}}$, are transferred into a target face, $I_{\text{target}}$, producing a composite image, $I_{\text{composite}}=f(I_{\text{source}},I_{\text{target}})$.
 (2) Facial Attribute Editing [18, 19], which alters specific attributes of a face, such as age, expression, or gender, expressed as $I^{\prime}=g(I,A_{\text{target}})$, where $A_{\text{target}}$ denotes the desired attribute.
 (3) Video-Driven Facial Animation [22, 23], which generates facial movements in static images using motion parameters $M$ from a driving source, represented by $I_{\text{animated}}=h(I,M)$.
-Category Type Methods # Videos Total
-Digital Manipulation Attribute-edit SAFA [27] 15,998 32,659 242,859
-StarGAN [28] 16,661
-Face-Swap FaceDancer [16] 16,168 176,878
-GHOST [17] 13,135
-SimSwap [18] 10,563
-InsightFace [19] 13,264
-BlendFace [20] 9,645
-DeepLiveCam [31] 12,059
-REFace [21] 12,756
-FOMM [22] 16,661
-InfoSwap [23] 16,534
-MegaFS [24] 14,334
-HifiFace [25] 16,661
-DiffSwap [26] 13,761
-Roop [32] 11,337
-Video-Driven DaGAN [29] 16,661 33,322
-OneShotTH [30] 16,661
-TABLE IV: The detailed statistics of digital manipulation attacks.
 Face-Swap.
@@
 OneShotTH [23] uses a novel keypoint representation to decompose identity and motion information unsupervisedly, enabling high-quality talking-head video synthesis with reduced bandwidth and interactive head rotation for enhanced video conferencing experiences.
-III-D2 Digital Adversarial Attacks
+TABLE III: Sample statistics for each protocol on the UniAttackData+ dataset, including train, dev, and test sets, where each column represents the number of samples for live faces or fake types.
+‘#’ indicates the number of videos.
+Protocol Subset #Live Physical Attacks Digtal Manipulation Digital Adversarial Digital Generation #All
+#2D #3D
+#Attribute-
+Edit
+#Face-
+Swap
+#Video-
+Driven
+#Pixel-
+Level
+#Semantic-
+Level
+#ID-
+Consistent
+#Style-
+Transfer
+#Prompt-
+Based
+P1 train 9,125 7,500 20,475 12,330 88,439 16,661 91,636 41,653 23,161 9,096 24,600 355,676
+dev 3,650 3,000 8,190 6,532 35,376 6,664 36,654 16,661 9,264 3,638 9,840 139,469
+text 5,475 4,500 12,285 13,797 53,063 9,997 54,981 24,991 13,896 5,458 14,759 213,202
+P2 train 13,650 5,880 19,110 11,663 89,522 11,663 81,639 34,988 21,078 9,096 23,026 321,315
+dev 1,000 2,520 8,190 4,998 38,366 4,998 34,988 14,995 9,033 3,638 9,868 132,594
+text 3,600 6,600 13,650 15,998 48,990 16,661 49,983 33,322 16,210 5,458 16,301 226,773
+P3 train 13,650 10,500 - 22,861 123,815 23,325 128,290 58,314 - - - 380,755
+dev 1,000 4,500 - 9,798 53,063 9,997 54,981 24,991 - - - 158,330
+text 3,600 - 40,950 - - - - - 46,321 18,192 49,199 158,262
+3.4.2 Digital Adversarial Attacks
 UniAttackData+ includes $16$ attacks designed to mislead machine learning models through subtle, often imperceptible perturbations.
@@
 Pixel-level attacks rely on precise, gradient-based adjustments, while semantic-level attacks manipulate intermediate representations or latent spaces, increasing their transferability and effectiveness.
-Category Type Methods # Videos Total
-Digital Adversarial Pixel_level ALMA [33] 16,661 183,271 266,576
-AdvDrop [34] 16,661
-AI_FGTM [35] 16,661
-ILA-DA [36] 16,661
-AutoAttack [37] 16,661
-PGD [38] 16,661
-Jitter [40] 16,661
-NIFGSM [41] 16,661
-PIFGSMpp [42] 16,661
-Pixel [43] 16,661
-VMIFGSM [44] 16,661
-Semantic-Level Demiguise [45] 16,661 83,305
-SSAH [46] 16,661
-DeCoWA [39] 16,661
-PNA_PatchOut [47] 16,661
-TGR [48] 16,661
-TABLE V: The detailed statistics of digital adversarial attacks.
 Pixel-Level.
@@
 Jitter [27] proposes a novel loss function that improves attack efficiency and success rates by better exploring the perturbation space and addressing model confidence issues.
-NI-FGSM [41] leverages Nesterov accelerated gradients, and SIM, optimizing perturbations across scaled copies, to enhance the transferability of adversarial examples, achieving higher success rates on defense models.
+NI-FGSM [76] leverages Nesterov accelerated gradients and SIM, optimizing perturbations across scaled copies, to enhance the transferability of adversarial examples, achieving higher success rates on defense models.
 PIM++ [77] enhances adversarial transferability by using a patch-wise iterative method with an amplified step size, a projection kernel for gradient redistribution, and temperature adjustments to balance transferability and white-box attack performance.
@@
 PNA-PatchOut [29] combines Pay No Attention (PNA), skipping attention gradients, and PatchOut, optimizing random patch subsets, to enhance adversarial transferability across Vision Transformers (ViTs) and CNNs.
-III-D3 Digital Generation Attacks
-UniAttackData+ comprise $7$ types leveraging diffusion models to produce highly realistic synthetic content.
+3.4.3 Digital Generation Attacks
+UniAttackData+ comprises $7$ types leveraging diffusion models to produce highly realistic synthetic content.
 These attacks are divided into three categories: Identity-Consistent Generation, Style Transfer, and Prompt-Driven Generation.
@@
 Prompt-Driven Generation [82] utilizes textual or image prompts to guide synthesis, expressed as $I_{\text{prompted}}=G(P_{\text{text}},P_{\text{img}})$.
-ID-Consistant.
+ID-Consistent.
 ConsistentID [31] enhances personalized portrait generation with a multimodal facial prompt generator for fine-grained control and an ID-preservation network with facial attention localization, achieving superior ID consistency and precision from a single reference image.
@@
 IP-Adapter [82] introduces a decoupled cross-attention mechanism to enable lightweight image prompting for text-to-image diffusion models, achieving multimodal generation with only 22M parameters, while maintaining compatibility with text prompts, structural controls, and custom models.
-Category Type Methods # Videos Total
-Digital Generation ID-Consistent ConsistentID [49] 15,436 46,321 113,712
-PuLID [50] 16,210
-InstantID [51] 14,675
-Style Transfer PhotoMaker [52] 18,192 18,192
-Prompt Based IP-Adapter_kolors [53] 16,449 49,199
-IP-Adapter_sdplusface [53] 16,449
-IP-Adapter_sdpxlface [53] 16,301
-TABLE VI: The detailed statistics of digital generation attacks.
-III-E Protocol
-ID Partition Protocol $P1$.
-Simply merging existing PAD and FFD datasets can cause models to focus on ID noise rather than discriminative features.
-To address this, we leverage the ID consistency of UniAttackData+.
-In this protocol, we train the model on a selected subset of identities, using the remaining identities for validation and testing.
-Each ID contains all attack types.
+3.5 Protocol and Statistic
+Identity Division Protocol $P1$.
+Simply merging existing PAD and FFD datasets can lead models to focus on identity-related noise rather than truly discriminative features.
+To mitigate this issue, we leverage the ID consistency provided by UniAttackData+.
+In this protocol, the model is trained on a selected subset of identities, while the remaining identities are reserved for validation and testing.
+Each identity includes all types of attacks to ensure comprehensive evaluation.
 Cross-attack Methods Protocol $P2$.
-In real-world deployment scenarios, a myriad of attack methods may emerge.
-Therefore, models should not only fit to seen attack methods but must also be capable of handling unseen attack methods.
-This protocol is designed to evaluate the model’s generalization ability across different attack methods.
-We adopted all identities for training and randomly selected 50% of attack methods as the training set, 20% as the validation set, and 30% as the test set.
-For instance, in the case of digital attacks, we used the first seven types of digital editing face swap attacks as the training set, two types for validation, and four types for testing, in order to assess the model’s performance on unseen face swap attacks.
+In real-world deployment scenarios, a wide range of attack methods may emerge.
+Therefore, models must not only perform well on known attack types but also generalize effectively to unseen ones.
+This protocol is designed to evaluate a model’s generalization capability across diverse attack methods.
+All identities are used for training, while attack methods are split randomly: 50% for training, 20% for validation, and 30% for testing.
+For example, in the case of digital attacks, the first seven types of digital face, i.e., swap attacks, are used for training, two types for validation, and four types for testing, enabling us to assess the model’s ability to handle unseen face-swap attacks.
 Cross-attack types Protocol $P3$.
-In many cases, models are trained only on simpler attack types but are tested against more advanced and complex attack methods.
+In many cases, models are trained only on simpler attack types but are tested against more advanced and complex attack types.
 This protocol aims to assess the model’s generalization ability from simpler to more sophisticated attacks.
 We trained the model using all identities, with 70% of physical 2D attacks, digital editing, and adversarial attacks from each identity as the training set, and 30% as the validation set.
-Physical 3D attacks and digital generative attacks were used as the test set, to evaluate the model’s generalization ability on unseen, more advanced attack types.
-IV Hierarchical Prompt Tuning (HiPTune)
+Physical 3D attacks and digital generative attacks were used as the test set to evaluate the model’s generalization ability on unseen, more advanced attack types.
+The sample statistics for each protocol are shown in Tab.
+III.
+We have detailed the number of train, dev, and test sets for each protocol, and also listed the types of spoofing contained in the subsets and their quantities.
+4 Hierarchical Prompt Tuning (HiPTune)
 Figure 4: (a) The VP-Tree establishes a hierarchical prompt tree for all attack types, organized into three levels of grouping categories.
-It assigns a learnable prompt for each attack type and integrates the hierarchical prompts into the final output.
-(b) The APP employs a hierarchical prompt pruning mechanism to select the most promising prompt for each sample.
-(c) The DPI aims to aggregate all candidate prompts from different levels of VP-Tree and projects them into the text encoder for subsequent similarity computation.
-IV-A Preliminaries
-Vision-Language Models (VLM) like CLIP [54] utilize both text encoder $\mathcal{T}(\cdot)$ (like Transformer [55]) and image encoder $\mathcal{I}(\cdot)$ (like ViT [56]), where the input is the (text, image) pair, to leverages broader supervision.
+It assigns a learnable prompt for each attack type and integrates the hierarchical prompts into different levels of the encoder.
+(b) The DPI aims to aggregate all candidate prompts from different levels of the VP-Tree and projects them into the text encoder for subsequent similarity computation.
+(c) The APP employs an adaptive prompt pruning mechanism to select the most promising prompt for each sample.
+4.1 Preliminaries
+Vision-Language Models (VLMs) like CLIP [83] utilize both text encoder $\mathcal{T}(\cdot)$ (like Transformer [84]) and image encoder $\mathcal{I}(\cdot)$ (like ViT [85]), where the input is the image-text pair, providing a wide range of semantic information as supervision.
 The outputs of both encoders are then dot-multiplied to obtain the probability score $p$:
-$p(y=i\mid\bm{x})=\frac{\exp\left(\cos\left(\bm{w}_{\bm{i}},\bm{f}\right)/\tau\right)}{\sum_{j=1}^{K}\exp\left(\cos\left(\bm{w}_{\bm{j}},\bm{f}\right)/\tau\right)},$ (1)
-where $\left\{\bm{w}_{i}\right\}_{i=1}^{K}$ denotes the set of weight vectors generated by the text encoder, and $\bm{f}$ represents the features extracted by the image encoder.
+$p(y=i\mid\boldsymbol{x})=\frac{\exp\left(\cos\left(\boldsymbol{w}_{\boldsymbol{i}},\boldsymbol{f}\right)/\tau\right)}{\sum_{j=1}^{K}\exp\left(\cos\left(\boldsymbol{w}_{\boldsymbol{j}},\boldsymbol{f}\right)/\tau\right)},$ (1)
+where $\left\{\boldsymbol{w}_{i}\right\}_{i=1}^{K}$ denotes the set of weight vectors generated by the text encoder, and $\boldsymbol{f}$ represents the features extracted by the image encoder.
 $K$ is the number of classes, and $\tau$ refers to the temperature parameter learned by CLIP, $\cos(\cdot,\cdot)$ means the cosine similarity.
-Moreover, CoOp [57] explores prompt engineering in the VLM, constructing learnable prompts $\bm{t}=[\mathrm{V}]_{1}[\mathrm{~{}V}]_{2}\ldots[\mathrm{~{}V}]_{M}[\mathrm{CLASS}]$ and feeding to the fixed Text Encoder.
+Moreover, CoOp [86] explores prompt engineering in the CLIP, constructing learnable prompts $\boldsymbol{t}=[\mathrm{V}]_{1}[\mathrm{~{}V}]_{2}\ldots[\mathrm{~{}V}]_{M}[\mathrm{CLASS}]$ and feeding to the frozen Text Encoder.
 The $[\mathrm{V}]_{m}(m\in\{1,\ldots,M\})$ is the vectors with the same dimension as word embeddings, and $M$ is a hyperparameter specifying the number of context tokens.
 The prediction probability is further computed as follows:
-$p(y=i\mid\bm{x})=\frac{\exp\left(\cos\left(\mathcal{T}\left(\bm{t}_{i}\right),\bm{f}\right)/\tau\right)}{\sum_{j=1}^{K}\exp\left(\cos\left(\mathcal{T}\left(\bm{t}_{j}\right),\bm{f}\right)/\tau\right)}.$ (2)
+$p(y=i\mid\boldsymbol{x})=\frac{\exp\left(\cos\left(\mathcal{T}\left(\boldsymbol{t}_{i}\right),\boldsymbol{f}\right)/\tau\right)}{\sum_{j=1}^{K}\exp\left(\cos\left(\mathcal{T}\left(\boldsymbol{t}_{j}\right),\boldsymbol{f}\right)/\tau\right)}.$ (2)
 Visual Prompt Tuning (VPT) aims to fine-tune the pre-trained models with only a few learnable parameters.
-Given the input image tokens $\bm{x}_{i},i\in{1,...,l}$ with length $l$, the trainable vectors $\bm{P}=\left\{\bm{p}^{k}\in\mathbb{R}^{d}\mid k\in\mathbb{N},1\leq k\leq p\right\}$, where d is the dimension of the patch embeddings and $p$ is the number of prompts, are cat with the original tokens and fed into the fixed encoders $Input=cat[\bm{x}\cdot\bm{P}]$.
-IV-B Overview of HiPTune
+Given the input image tokens $\boldsymbol{x}_{i},i\in{1,...,l}$ with length $l$, the trainable vectors $\boldsymbol{P}=\left\{\boldsymbol{p}^{k}\in\mathbb{R}^{d}\mid k\in\mathbb{N},1\leq k\leq p\right\}$, where $d$ is the dimension of the patch embeddings and $p$ is the number of prompts, are cat with the original tokens and fed into the frozen encoders $Input=cat[\boldsymbol{x}\cdot\boldsymbol{P}]$ in certain orders.
+4.2 Overview of HiPTune
 We proposed a HiPTune that adopts hierarchical prompts to find multiple classification criteria from different semantic spaces and further integrates them, aiming to form more robust classification criteria for each input sample.
 Shown in Fig.
-3, first, we construct the Visual Prompt Tree to gain the various semantics spaces tailored for each attack detection task.
-Then, given the input image $\bm{x}$, we adaptively select the most suitable prompt from each level of the VP-Tree by our Adaptive Prompt Pruning module.
-In addition, we integrate the prompt candidates from each selected level of the APP to describe the complete features of the input sample and feed it to the text encoder for better alignment.
-IV-C Visual Prompt Tree (VP-Tree)
+4, first, we construct the Visual Prompt Tree (VP-Tree) to gain the various semantics spaces tailored for each attack detection task.
+Then, given the input image $\boldsymbol{x}$, we adaptively select the most suitable prompt from each level of the VP-Tree by our Adaptive Prompt Pruning (APP) module.
+In addition, we integrate the prompt candidates from each selected level of the APP to describe the complete features of the input sample and feed it to the text encoder for better alignment through our Dynamic Prompt Integration (DPI) mechanism.
+4.3 Visual Prompt Tree (VP-Tree)
+Inspired by the work of [87, 88], learnable prompt parameters can be used to adapt a pretrained encoder, steering its output features to better align with downstream tasks.
+In the context of the UAD task, the diversity and high realism of forgery techniques pose significant challenges.
+To address this, we propose a hierarchical prompt tree designed to deliver coarse-to-fine guidance throughout the layered processing of a Transformer.
+At each block, the prompts encourage the encoder to refine its feature representations progressively, improving discrimination for classification.
+Notably, our prompts are class-agnostic in the sense that they target the distinction between authentic and forged samples in general, applying a specific prompt to each “real–fake” sample pair to capture their distinguishing characteristics.
 Define the levels.
-We can obtain a hierarchical relationship between each attack type based on the spoof and forgery types.
+We can obtain a hierarchical relationship between each attack type based on the spoof and forgery types from Fig.
+4.
 Apart from the subjective categorization from the dataset, the features of different attacks demonstrate a more compact grouping relative to other categories while maintaining distinct independence within each category.
@@
 4 (a).
-The first level distributes the most obvious data as live, physical, and digital attack subtrees.
-The physical subtree is first separated into 2D and 3D attacks, which represents the second level of the VP-Tree.
+The first level distributes the most obvious data as physical and digital attack subtrees.
+The physical subtree is first separated into 2D and 3D attacks, which represent the second level of the VP-Tree.
 Print, replay, and cutouts are three key types for the level 3 prompts of a 2D attack.
@@
 Concerning the digital subtree, we categorize the digital attacks into $3$ kinds according to their forgery techniques.
-The digital manipulation method includes attribute-editing, face-swap, and video-driven faces.
+The digital manipulation method includes attribute editing, face swapping, and video-driven faces.
 Adversarial attacks are divided into pixel-level and semantic-level attacks for forgery areas.
 We also define $3$ sub-prompts for the digital generation attacks: ID-consistent generation, style transfer, and prompt-based editing forgeries.
+It is worth noting that to ensure continuity in the properties and semantics at each level, each subsequent prompt partially inherits semantic information from its preceding prompt.
+Specifically, given a prompt defined in $\boldsymbol{p}^{0}\in\mathbb{R}^{l\times D}$, we apply an inheritance coefficient $\alpha$ to generate the inherited prompt $\boldsymbol{p}^{1}=cat[\boldsymbol{p}^{1}\in\mathbb{R}^{(l-\alpha l)\times D},\boldsymbol{p}^{1}\in\mathbb{R}^{\alpha l\times D}]$.
+This inherited component is then proportionally compressed, every $\frac{\alpha}{1}\cdot l$ tokens have been calculated the mean value and concatenated to the next-level prompt.
+This design ensures that the new prompt retains essential characteristics from its parent while introducing its own fine-grained, task-specific distinctions.
 Initialize the prompts. After we obtain the VP-Tree’s structure, we define the prompts for each level with random initialization in the following form:
-$\bm{p}=[\mathrm{V}]_{1}[\mathrm{~{}V}]_{2}\ldots[\mathrm{~{}V}]_{l},$ (3)
+$\boldsymbol{p}=[\mathrm{V}]_{1}[\mathrm{~{}V}]_{2}\ldots[\mathrm{~{}V}]_{l},$ (3)
 where each $[\mathrm{V}]_{l}(l\in\{1,\ldots,L\})$ is the learnable vectors with the length of $L$ at the same dimension with the patch embeddings (i.e., 768 for CLIP image encoder).
@@
 In order to force each prompt to learn the classification criteria specific to each attack type, we train the prompts in a supervised manner during the first stage.
-Given the input $\bm{x}\in\mathbb{R}^{C\times H\times W}$, where $C,H,W$ represent the channel, height, and width of the input images, we manually record the attack types and select the corresponding prompts at each level accordingly.
-For the live face images, since we train our model with (live, fake) pairs, we choose the same prompt as the partner.
-Then, we average the selected prompts to form the final prompt and link with the image tokens.
-The final input of the image encoder is formulated as Eq.
+Given the input $\boldsymbol{x}\in\mathbb{R}^{C\times H\times W}$, where $C,H,W$ represent the channel, height, and width of the input images, we manually record the attack types and select the corresponding prompts at each level accordingly.
+Following the approach of XX, we insert the prompt tokens between the original image tokens and the image class token.
+This design allows the prompts to effectively modulate the flow of information from the input representations to the classification token, encouraging the encoder to integrate the prompt’s guidance during feature aggregation and thereby refining task-relevant representations.
+The final input of each Transformer block is formulated as Eq.
 4:
-${\mathcal{I}}_{input}=cat[(\bm{x}_{cls},\bm{x}_{1},...,\bm{x}_{l})\cdot\bm{p}],$ (4)
-where $(\bm{x}_{cls},\bm{x}_{1},...,\bm{x}_{l})\in\mathbb{R}^{l+1,dim}$ is the image tokens after the patch embeddings.
+${\mathcal{I}}_{input}=cat[\boldsymbol{x}_{cls},\boldsymbol{p},(\boldsymbol{x}_{1},...,\boldsymbol{x}_{l})]$ (4)
+where $(\boldsymbol{x}_{cls},\boldsymbol{x}_{1},...,\boldsymbol{x}_{l})\in\mathbb{R}^{l+1,dim}$ is the image tokens after the patch embeddings.
 $l$ represents the length of image tokens, and $dim=768$ is the embedding dimension.
-The $\bm{p}$ is the final output of the integrated prompts as Eq.
-5
-$\bm{p}=mean(\bm{p}_{1}+\bm{p}_{2}+\bm{p}_{3}),$ (5)
-where $\bm{p}_{1},\bm{p}_{2},\bm{p}_{3}$ are the selected prompts at level $1,2,3$ and $mean(\cdot)$ represents the average of the inputs.
-After fully training the VP-Tree to represent the classification criteria for multiple semantic spaces, the Adaptive Prompt Pruning (APP) module will learn to select the most suitable prompts for each sample.
-IV-D Adaptive Prompt Pruning (APP)
-As shown Fig.
-3 (b), aiming to explore classification criteria from different semantic spaces, the APP module adopts a hierarchical pruning mechanism to select the most suitable prompt for each sample.
-Given the input image $\bm{x}\in\mathbb{R}^{C\times H\times W}$, the patch embedding first projects the image into the visual embedding dimension and reshapes $\bm{x}$ into a series of image token with length $L$: $\bm{x}:\mathbb{R}^{C\times H\times W}\rightarrow\mathbb{R}^{L\times D}$.
-In our case, $C=3$ represents the channels for the image, and $H,W$ are both $244$ for the input image, the token’s length $L=196$, and the dimension of visual embedding is $D=768$.
-After adding the positional embeddings, a class token is linked with the image token, and then the image token $\bm{x}\in\mathbb{R}^{197\times 768}$ is fed to the APP.
-Level-1 prompt pruning.
-Given the significant distribution variance between physical and digital attacks, it is simple to determine whether physical or digital criteria should be used to classify this sample.
-Thus, we utilize a simple linear layer to select the level 1 prompt:
-$G_{1}(\bm{x})=\mathbf{W}_{1}\cdot\bm{x}+bias_{1},$ (6)
-where $G_{1}$ is the gate function of the level-1 prompt pruning and $\mathbf{W}_{1},bias_{1}$ are learnable weights and bias.
-The level-1 router first projects each patch’s features into three alternatives using $G_{1}$ and then calculates the mean value over all tokens.
-After applying Softmax between three classes, we choose the $TOP_{K=1}$ candidate prompt.
-The whole procedure can be described as:
-$i^{p}_{1}=TOP_{k=1}(softmax(mean(G_{1}(\bm{x}),197))),$ (7)
-where $i^{p}_{1}$ denotes the index of the selected prompt, $mean(\cdot,\cdot)$ represents the mean calculus over specific dimension.
-Furthermore, the selected prompt is linked with the original image tokens and forwarded to the next level.
-Level-2 prompt pruning.
-Regarding the level-2 prompt pruning, it is challenging to separate samples from attack types within each category if we only rely on fixed visual embeddings.
-In light of this, we apply Central Difference Convolution (CDC) [58] and the Frequency-Adaptive Dilated Convolution (FADC) [59] for level-2 prompt pruning in the physical and digital branches, respectively.
-After the CDC or the FADC, the output feature map is sent to a linear layer, and then the tokens’ scores are averaged to select the candidate prompt further.
-The level-2 gate functions can be expressed as:
-$G_{2}^{0}$ $=\mathbf{W}_{2}^{0}\cdot CDC_{2}^{0}(\bm{x})+bias_{2}^{0}$ (8)
-$G_{2}^{1}$ $=\mathbf{W}_{2}^{1}\cdot FADC_{2}^{1}(\bm{x})+bias_{2}^{1},$
-and the selected prompt can be denoted as:
-$i^{p}_{2}=TOP_{k=1}(softmax(mean(G_{2}(\bm{x}),197))).$ (9)
-The level-1 and level-2 prompts are combined averagely and concat with the image tokens and fed to the next router.
-Level-3 prompt pruning. The level-3 prompt pruning procedure is the same as the level-2 pruning process. The gate functions can be described as follows:
-$G_{3}^{i}=\mathbf{W}_{3}^{i}\cdot Conv_{3}^{i}(\bm{x})+bias_{3}^{i},$ (10)
-where $i\in{1,2,3}$ for physical 2D/3D, digital manipulation, generation, and $i\in{1,2}$ for digital adversarial attacks.
-Finally, the selected prompts from all three levels are averaged together and concated with the original image tokens.
-For the live images, the live prompt from the first level will be directly concated with the tokens and fed to the image encoder.
-IV-E Dynamic Prompt Interaction (DPI)
+The $\boldsymbol{p}$ is the selected prompt for each level.
+As illustrated in the figure, we integrate different levels of prompts into specific layers of the Transformer blocks.
+As the Transformer progressively deepens, the semantic representation of tokens becomes increasingly precise.
+To leverage this property, we insert the first-level prompt after $N_{0}$ blocks—once the features have developed sufficient basic representational capacity—focusing on coarse discrimination between physical and digital attacks.
+Subsequently, after an additional $N_{1}$ and $N_{2}$ blocks, we introduce the second and third-level prompts, which provide increasingly fine-grained semantic guidance.
+Finally, the features undergo further processing through $N_{3}$ additional blocks to yield the final representation for classification.
+After fully training, the VP-Tree to represent the classification criteria for multiple semantic spaces, the Adaptive Prompt Pruning (APP) module will learn to select the most suitable prompts for each sample.
+TABLE IV: Results on UniAttackData+. CLIP-V [83] indicates removing the text encoder from CLIP. CoOp [86] (unified/specific) represents the CoOp without/with the class-specific context.
+Model P1: ID P2: cross-attack-methods P3.: cross-attacks-types
+ACER AUC ACC EER ACER AUC ACC EER ACER AUC ACC EER
+CLIP-V [83] 9.36 92.46 88.05 9.36 28.77 71.99 78.49 28.77 49.87 43.12 58.08 49.87
+CLIP [83] 9.93 95.22 83.75 9.93 22.89 72.79 76.15 22.89 34.37 58.09 59.90 34.37
+CoOp [86] 14.76 90.89 83.92 14.76 43.14 59.64 47.83 43.14 72.00 20.89 27.18 72.00
+CDCN++ [89] 12.77 94.53 90.89 11.72 15.23 91.78 84.72 15.22 37.55 67.90 58.67 38.46
+CFPL-FAS [90] 13.93 93.77 96.24 11.65 16.10 86.34 83.09 16.10 29.60 73.50 71.46 29.60
+MoE-FFD [91] 17.30 91.90 96.50 15.69 30.53 70.09 92.82 31.39 46.01 51.27 89.14 47.97
+LLA-Net [92] 10.57 95.64 95.97 7.61 18.34 82.97 89.86 21.31 36.58 57.02 68.25 42.20
+FA3-CLIP [59] 8.26 93.96 94.53 8.45 35.81 63.89 88.51 47.37 44.44 48.34 56.95 41.05
+HiPTune 7.31 97.87 90.02 7.61 14.26 92.50 90.26 14.71 25.68 77.62 69.34 26.02
+4.4 Adaptive Prompt Pruning (APP)
+As shown in Fig.
+4 (a) and (c), aiming to explore classification criteria from different semantic spaces, the APP module adopts a hierarchical pruning mechanism to select the most suitable prompt for each sample.
+Given the input image $\boldsymbol{x}\in\mathbb{R}^{C\times H\times W}$, the patch embedding first projects the image into the visual embedding dimension and reshapes $\boldsymbol{x}$ into a series of image tokens with length $L$: $\boldsymbol{x}:\mathbb{R}^{C\times H\times W}\rightarrow\mathbb{R}^{L\times D}$.
+In our case, $C=3$ represents the channels for the image, and $H,W$ are both $224$ for the input image, the token’s length $L=196$, and the dimension of visual embedding is $D=768$.
+After adding the positional embeddings, a class token is linked with the image token, and then the image token $\boldsymbol{x}\in\mathbb{R}^{197\times 768}$ is fed to the Transformer blocks.
+Through $N_{0}$ blocks, the image tokens and the level-1 prompts are fed into the APP.
+To select the most appropriate prompt, we first designate the front-most element (the first token after concatenation, with length 1) as the representative of each prompt.
+For the input image tokens, we use their $[CLS]$ token as the representation of the entire image.
+We then compute the attention scores between all prompt representatives and the image $[CLS]$ token.
+After applying the softmax function to normalize these scores, we select the single most relevant prompt using a $TOP_{k=1}$ strategy.
+This selected prompt is then inserted back into the sequence of image tokens to guide further encoding.
+The overall procedure is formally described as follows:
+$index=TOP_{k=1}[Attention(\textbf{p}[0,:]\cdot x[0,:,:])],$ (5)
+where $Attention(\cdot)$ represents the attention calculation, and the $index$ is the selection index of the most suitable prompt.
+It is important to note that to ensure each prompt accurately captures its intended semantic information during selection, we supervise the training of prompts using ground-truth labels at each level for different samples in the training set.
+This enables each prompt to specialize in discriminating its designated category or property.
+Once trained, the resulting VP-Tree is integrated into the model as a fixed module, supporting the APP (Adaptive Prompt Pooling) selection mechanism during downstream training.
+4.5 Dynamic Prompt Interaction (DPI)
 CLIP’s default templates,  e.g., “This is a photo of [CLASS]”, fail to describe the detailed semantic information for the multiple attack types, which caused the performance drops encountering advanced attacks.
@@
 Thus, we proposed a Dynamic Prompt Interaction (DPI) to project the visual prompts to the text encoder to fully describe the abundant features corresponding to each category.
-Specifically, first, we directly pass the live prompts from the first level as the live class text prompts.
-As for the fake prompts, we apply the probabilities that the app has predicted to the candidate prompts and combine them weightedly.
-For the first level, we apply the softmax function only to physical and digital prompt scores to obtain the binary probabilities of two categories.
-Then, we weighted sum the prompts to get a weighted level-1 fake prompt.
-In the second level, due to the router from the first level choosing a subtree for the sample and needing a possible fake prompt specific to the input sample, we repeat the weighted sum of candidate prompts until the router has not predicted the further probabilities.
-Lastly, we form prompts for both live and fake faces and send them to the text encoder for further similarity computation.
-IV-F Objectives and Inference
-The training objective for the first training stage is to help the prompts to learn the semantic information and classification criteria specific to the attack type.
-The second stage aims to train the APP to select the most suitable prompts tailored for each sample successfully.
-The loss for both stages is:
-$\mathcal{L}=\mathcal{L}_{CE}+\mathcal{L}_{triple},$ (11)
-where $\mathcal{L}_{CE}$ represents the cross-entropy loss and $\mathcal{L}_{triple}$ is the Asymmetric Triplet Loss [60].
-The reason for applying Asymmetric Triplet Loss is to separate the different prompts to preserve their unique classification criteria while combat the relatively similar prompts.
-V Experiments
-V-A Experiment Settings
+Specifically, shown as Fig.
+4 (b), after the APP module selects the appropriate prompts at each hierarchical level, we stack these prompts in order to form a combined sequence.
+This stacked sequence is then processed using self-attention, allowing information exchange across levels.
+Following the softmax operation, we compute a weighted sum of the three hierarchical prompts to produce the final unified prompt.
+Finally, this prompt is projected from the original image feature dimension to the target text feature dimension through a dedicated mapping layer.
+5 Experiments
+5.1 Experiment Settings
 Datasets, Protocols, and Evaluation Metrics.
-We evaluated the effectiveness of the proposed UniAttackData+ dataset and the robustness of the HiPTune algorithm using three existing UAD datasets.
-For JFSFDB, we conducted both internal and external testing within its original FAS and FFD datasets, as well as within the joint training dataset.
-For UniAttackData, we tested using its original p1, p2.1, and p2.2 subsets.
-For our proposed UniAttackData+, we performed training and testing based on the four protocols we introduced.
-Additionally, we established a UAD-Benchmark to assess the model’s performance in cross-dataset testing.
-We evaluate model performance with four metrics: (1) ACER (average of FRR and FAR), (2) AUC (theoretical model performance), (3) EER (optimal threshold), and (4) ACC (overall accuracy).
+We evaluate the effectiveness of the proposed UniAttackData+ and HiPTune using three existing UAD datasets.
+For the UniAttackData+ dataset, we conduct experiments on the three proposed protocols.
+For JFSFDB [38] dataset, we only conduct cross-domain experiments on both separate training on each single task (i.e., PAD or FFD) and joint training within both two tasks (i.e., UAD).
+Specifically, separate training is conducted by training two distinct models on datasets SiW, 3DMAD, HKBU for PAD, and FF++ for FFD, respectively.
+The trained models are then evaluated on cross-domain datasets MSU, 3DMask, ROSE for PAD and DFDC, CelebDFv2 for FFD.
+Instead of training separate models, joint training involves learning a unified model for both tasks by leveraging all data from SiW, 3DMAD, HKBU, FF++.
+The unified model is then evaluated on cross-domain datasets MSU, 3DMask, ROSE, DFDC, CelebDFv2 for UAD.
+For UniAttackData [39] dataset: (1) Protocol 1 evaluates performance under a unified attack detection task with all attack types included in training and testing; (2) Protocol 2 assesses generalization to unseen attack types using a leave-one-type-out strategy, with Protocols 2.1 and 2.2 focusing on unseen physical and digital attacks, respectively.
+Following the setup in FA3-CLIP [59], we also conduct experiments under Protocol 1.1, 1.2, and 1.3.
+Protocols 1.1 and 1.2 independently evaluate the digital category by excluding deepfake and adversarial attacks, respectively, from training and validation, testing them only on separate identity sets.
+Protocol 1.3 includes all digital subtypes with standard distribution.
+These protocols assess the model’s robustness in unified face attack detection across both physical and digital threats.
+In addition, we construct a domain generalization-based UAD benchmark using the JFSFDB, UniAttackData, and UniAttackData+ datasets, where any two are combined for training and the remaining one is used for testing.
+We evaluate performance using four metrics: (1) ACER, the average of the false rejection rate (FRR) and the false acceptance rate (FAR), (2) AUC, reflecting the model’s theoretical performance, (3) EER, representing the error rate at the optimal threshold, and (4) ACC, indicating the overall classification accuracy.
 Implementation Details.
-We use a $3$-level VP-Tree with a prompt length of $60$, a $\theta$-value of CDC set to $0.7$, and a fixed CLIP model with ViT/b-16 backbone, classifying ”live” and ”fake.” The initial learning rate is $10^{-5}$ with the Adam optimizer.
+We use a $3$-level VP-Tree with a prompt length of $40$ and a frozen CLIP model with ViT/b-16 backbone, classifying “live” and “fake”.
+The initial learning rate is $10^{-5}$ with the Adam optimizer.
 Images are resized to $224\times 224$ with random cropping and flipping.
-VP-Tree training runs on an NVIDIA GeForce RTX 4090 for $100$ epochs (batch size: $32$), followed by full model training for $200$ epochs (batch size: $32$).
-V-B Results on UniAttackData+
-p1-ID p2-cross_attack_methods p3.1-cross_attacks(low2high) p3.2-cross_attacks(high2low)
-Model ACER AUC ACC EER ACER AUC ACC EER ACER AUC ACC EER ACER AUC ACC EER
-CLIP-V 9.36 92.46 88.05 9.36 28.77 71.99 78.49 28.77 49.87 43.12 58.08 49.87 51.17 47.35 86.76 51.17
-CLIP 9.93 95.22 83.75 9.93 22.89 72.79 76.15 22.89 34.37 58.09 59.9 34.37 53.05 46.95 82.68 53.05
-CoOp(unified) 14.76 90.89 83.92 14.76 43.14 59.64 47.83 43.14 72 20.89 27.18 72 45.06 53.7 59.2 45.06
-CoOp(specific) 15.26 90.8 84.48 15.26 44.73 55.44 48.5 44.73 71.57 21.76 24.79 71.57 42.82 57.26 57.83 42.82
-CDCN++ 12.77 94.53 90.89 11.72 15.23 91.78 84.72 15.22 37.55 67.9 58.67 38.46 42.09 62.16 70.55 39.66
-CFPL-FAS 13.93 93.77 96.235 11.65 16.1 86.34 83.09 16.1 29.6 73.5 71.46 29.6 53.2 48 84.98 50.21
-MoE-FFD 17.3 91.9 96.5 15.69 30.53 70.09 92.82 31.39 46.01 51.27 89.14 47.97 54.85 45.11 88.81 54.85
-LLA-Net 10.57 95.64 95.97 7.61 18.34 82.97 89.86 21.31 36.58 57.02 68.25 42.2 54.04 45.94 87.71 54.04
-FA-CLIP 8.26 93.96 94.53 8.45 35.81 63.89 88.51 47.37 44.44 48.34 56.95 41.05 51.22 48.78 90.38 99.64
-HiPTune 7.31 97.87 90.02 7.61 14.26 92.5 90.26 14.71 25.68 77.62 69.34 26.02 40.13 65.89 92.66 40.18
-TABLE VII: Results on UniAttackData+. CoOp(unified/specific) represents the CoOp without/with the class-specific context.
-We employed a range of CLIP-based baseline models (including CLIP-V, CLIP, CoOp without class-specific context, and CoOp with class-specific context) to evaluate the performance of shared-space learning models on our dataset.
-In addition, we assessed algorithms specifically designed for PAD/FFD, such as CFPL-FAS, MoE-FFD, and LLA-Net, to determine their generalization capabilities.
-We also compared our approach with the recently released $FA^{3}-CLIP$, an algorithm specifically designed for unified attack detection.
-The detailed experimental results are presented in Table VII.
-Authenticity-Only Testing Protocol $p1$:
-This protocol primarily aims to evaluate the model’s true capability in distinguishing between real and fake samples.
-As shown in the experimental results table (Table VII), models that attempt to update the encoder generally achieve better performance.
-This is because, within UniAttackData+, a shared feature space still exists where different attack types can be clustered, and enhancing the image encoder’s feature extraction capability improves model performance.
-However, these models have a large number of parameters, making them unsuitable for deployment in real-world environments.
-Conversely, CoOp, which relies on prompt matching, attempts to map fixed features to relevant regions.
-However, given the vast and unpredictable variety of attack types, a single prompt is insufficient to address the diversity and sophistication of these attacks.
-Models specifically designed for PAD and FFD focus solely on the manipulated regions of the respective images, neglecting the common forgery cues and underlying processes shared across UAD tasks.
-Notably, CDCN++, a purely convolutional model, also demonstrates decent performance, highlighting the importance of local region analysis in UAD tasks.
-This suggests that local feature extraction remains crucial for robust attack detection.
-Our HiPTune adopts a visual prompt approach, which not only guides the encoder in feature extraction but also maintains model efficiency.
-Furthermore, our hierarchical prompt tree encapsulates multi-level, fine-grained classification criteria, achieving highly precise classification performance.
-Real-World Robustness Testing Protocol $p2$:
-This protocol is widely applicable in real-world scenarios, aiming to evaluate the model’s performance on seen attack types (such as physical 2D attacks) and unseen attack methods (such as face-swapping techniques like Roop and SimSwap).
-Similar to protocol p1, models that train the encoder generally achieve superior performance.
-This is because identifying a shared feature space helps the model generate a general mapping for previously unseen methods, effectively leveraging existing knowledge to handle new attack techniques.
-In contrast, models like CoOp, which rely on prompt matching, are more prone to overfitting the training data.
-As a result, their detection performance significantly declines when encountering new data distributions.
-For instance, CDCN++, a model specifically designed for PAD, utilizes a central difference convolutional structure that emphasizes local detail information, achieving the best results in our benchmark experiments.
-This finding highlights the importance of focusing on local detail information to ensure model generalization in UAD tasks.
-Our HiPTune model adopts a visual prompt approach, introducing prompts on the image encoder side to guide the frozen encoder in extracting relevant features while maintaining efficiency.
-Unlike conventional prompt-matching methods, our prompts serve only as guidance, avoiding excessive dependence on existing knowledge for decision-making.
-This design not only enhances the model’s generalization capability against unseen attack methods but also ensures robustness and interpretability.
-Upward Compatibility Protocol $p3.1$:
-This protocol aims to assess whether models trained on simple attack types can maintain high robustness in testing environments featuring complex and advanced attacks.
-We observe a sharp performance decline in both CoOp models, which we attribute to the absence of encoder guidance.
-Relying solely on text-side mapping relationships fails to achieve strong reasoning and generalization capabilities.
-In contrast, our HiPTune model, which adaptively integrates different types of prompts, demonstrates superior performance.
-Despite being trained on limited samples, HiPTune can make judgments based on existing higher-level prompts while leveraging data from other attack types to assist in inference.
-This design ensures that the model retains strong generalization even in the presence of advanced attack types.
-Downward Compatibility Protocol $p3.2$:
-This protocol aims to evaluate whether models can leverage advanced knowledge to accurately detect simpler attack types.
-As shown in the table, most models exhibit poor performance, indicating that existing methods are insufficient to effectively utilize anti-forgery knowledge derived from advanced digital generative attacks.
-This also reveals the potential non-transferability between these two types of knowledge.
-In contrast, our HiPTune model demonstrates superior performance.
-It effectively borrows knowledge from higher-level prompts and dynamically combines existing classification criteria, highlighting its advanced design.
-This adaptive approach allows HiPTune to generalize knowledge from complex attacks to simpler ones, ensuring robust performance across diverse scenarios.
-V-C Results on UAD-Benchmark
-JFSFDB-UniAttackData+ UniAttackData-UniAttackData+ JFSFDB+UniAttackData-UniAttacData+
-ACER AUC ACC EER ACER AUC ACC EER ACER AUC ACC EER
-CLIP-V 35.83 67.91 53.44 35.83 41.41 60.9 49.69 41.41 39.81 63.4 54.1 39.81
-CLIP 39.32 63.95 55.65 39.32 40.58 62.75 50.38 40.58 39.36 64.08 51.92 39.36
-CoOp(unified) 39.59 63.56 58.29 39.6 26.93 78.77 70.44 26.93 29.77 76.77 67.4 29.77
-CoOp(specific) 39.39 63.89 58.71 39.39 28.62 77.5 69.9 28.62 30.3 75.9 68.49 30.3
-HiPTune 32.88 69.34 60.51 32.88 20.16 88.53 86.82 20.23 25.93 83.05 73.82 25.94
-UniAttackData-JFSFDB UniAttackData+-JFSFDB UniAttackData+UniAttacData+-JFSFDB
-ACER AUC ACC EER ACER AUC ACC EER ACER AUC ACC EER
-CLIP-V 36.16 64.97 70.81 36.16 51.65 48.75 47.88 51.65 45.12 56.26 50.52 45.12
-CLIP 36.98 64.05 76.04 36.98 50.72 48.25 44.09 50.72 44.42 58.1 54.11 44.42
-CoOp(unified) 40.02 64.12 58.53 40.02 43.55 59.51 54.28 43.55 38.21 65.84 59.44 38.21
-CoOp(specific) 39.44 64.83 59.02 39.44 42.95 59.99 55.39 42.95 38.13 66.04 60.67 38.13
-HiPTune 34.39 71.23 80.31 34.42 28.14 79.22 70.46 28.78 32.93 70.89 72.23 32.92
-JFSFDB-UniAttackData UniAttackData+-UniAttackData JFSFDB+UniAttackData+-UniAttacData
-ACER AUC ACC EER ACER AUC ACC EER ACER AUC ACC EER
-CLIP-V 29.94 75.58 62.11 29.94 22.26 84.22 56.92 22.26 13.15 91.86 77.74 13.15
-CLIP 29.27 80.12 67.63 29.27 22.38 84.11 75.76 22.38 16.59 90.14 82.5 16.59
-CoOp(unified) 37.17 67.2 58.91 37.17 25.4 80.5 71.77 25.4 30.28 76.11 65.41 30.28
-CoOp(specific) 41.4 61.95 57.39 41.4 26.37 79.32 70.68 26.37 27.34 79.19 67.3 27.34
-HiPTune 21.36 85.59 74.39 31.56 18.62 90.75 88.64 18.63 8.55 94.67 90.26 8.56
-TABLE VIII: Results on existing UAD datasets.
-In the UAD-Benchmark, training and testing models across different datasets allow us to evaluate not only their generalization to various attack types but also their domain generalization capabilities.
-By examining the table, we can draw two key insights:
-(1) As shown in the second column of the table, models trained on UniAttackData+ demonstrate strong anti-forgery capabilities that can be effectively transferred to other datasets.
-This highlights the robustness and generalizability of the knowledge learned from UniAttackData+.
-(2) Baseline models show limited improvement, suggesting that we should design models specifically tailored to the characteristics of the dataset.
-Such models should be capable of fully leveraging the knowledge contained within the dataset, maximizing its utility for model training and evaluation.
-V-D Results on JFSFDB
-Face Spoofing Detection Face Forgery Detection
-Intra-testing Cross-testing Intra-testing Cross-testing
-Metrics Sep./ Joint Method SiW 3DMAD HKBU MSU 3DMask ROSE FF++ DFDC Celeb-DFv2 Average
-MesoNet 95.73 100 68.87 85.33 44.54 68.05 65.54 54.72 65.1 71.99
-Xception 99.88 100 94.47 90.62 63.88 90.15 98.21 64.86 84.27 87.37
-MultiAtten 99.56 100 68.68 89.21 36.88 72.66 70.24 63.37 64.34 73.88
-CDCN++ 100 99.56 94.51 82.65 47.66 76.76 97.53 56.65 78.34 81.51
-DeepPixel 100 100 89.4 95.17 56.87 80.6 99.26 67.54 85.51 86.04
-ResNet50 99.99 100 93.71 94.33 58.95 88.79 97.04 62.35 82.51 86.41
-ViT 100 100 95.45 95.13 35.6 89.12 98.65 73.79 86.61 86.04
-CLIP-V 99.99 100 96.85 97.29 76.62 93.6 93.05 82.09 75.4 90.54
-CLIP 99.99 100 95.58 95.82 80.4 92.89 95.92 86.08 79.4 91.79
-CoOp(Unified) 99.86 100 94.42 83.73 91.94 94.63 79.49 64.46 62.05 85.62
-CoOp(Specific) 99.83 100 93.72 84.14 92.61 92.89 81.41 66.63 61.88 85.9
-Sep. HipTune 99.97 100 98.62 98.62 95.87 98/62 100 89.72 89.63 93.12
-MesoNet 92.55 97.94 78.42 84.01 45.87 70 56.66 50.57 55.03 72.56
-Xception 99.8 100 97.62 91.56 75.49 79.38 96.69 63.69 78.87 87.01
-MultiAtten 99.38 100 78.34 92.89 58.31 77.51 69 58.11 60 77.06
-CDCN++ 99.99 98.89 96.8 78.93 61.54 76.39 95.79 56.42 76.92 82.41
-DeepPixel 100 100 97.31 93.97 66.65 74.58 99.22 66.36 88.45 87.39
-ResNet50 99.93 100 98.1 93.5 79.64 86.03 98.83 67.18 85.55 89.86
-ViT 100 100 100 95.85 62.74 91.88 98.83 75.11 88.36 90.31
-JFSFDB-Uni 100 100 99.96 97.99 67.1 95.57 99.12 75.9 89.76 91.71
-CLIP-V 100 100 95.91 97.23 71.26 83.96 95.69 82.84 79.64 89.44
-CLIP 100 100 97.48 96.39 86.13 81.07 94.83 84.06 78.99 90.83
-CoOp(Unified) 99.71 100 95.82 81.58 89.27 93.62 78.55 68.56 64.88 85.78
-CoOp(Specific) 99.78 100 95.33 80.68 92.82 93.34 76.42 67.58 65.06 85.67
-AUC(%) Joint HipTune 100 100 100 99.62 95.43 97.82 100 90.88 93.35 97.46
-TABLE IX: Results on JFSFDB AUC.
-Face Spoofing Detection Face Forgery Detection
-Intra-testing Cross-testing Intra-testing Cross-testing
-Metrics Sep./ Joint Method SiW 3DMAD HKBU MSU 3DMask ROSE FF++ DFDC Celeb-DFv2 Average
-MesoNet 11.67 0 35.12 21.9 55.82 36.81 38.93 46.16 38.78 31.69
-Xception 1.14 0 11.31 18.57 42.3 18.13 5.54 38.92 24.07 17.78
-MultiAtten 2.78 0 35.71 18.57 61.22 33.82 35.54 39.73 40.46 29.76
-CDCN++ 0.14 3.33 11.9 22.8 52.76 32.01 7.5 44.75 28.19 22.6
-DeepPixel 0 0 17.86 15.71 46.42 28.23 2.32 36.82 22.81 18.91
-ResNet50 0.36 0 13.69 12.38 45.24 19.94 8.18 40.61 26.01 18.49
-ViT 0.14 0 11.31 11.9 60.4 19.47 4.64 32.75 22.76 18.15
-CLIP-V 0.49 0 7.74 8.58 29.56 13.17 13.66 25.43 29.89 14.28
-CLIP 0.34 0 11.51 12.86 26.74 14.61 11.47 23.14 27.16 14.2
-CoOp(Unified) 1.51 0 12.7 25.95 17.02 12.06 27.14 39.42 41.21 19.67
-CoOp(Specific) 1.66 0 15.08 25.48 15.68 15.25 25.98 37.69 41.53 19.82
-Sep. HipTune 0.35 0 4.58 6.56 11.57 12.82 1.89 20.86 28.77 9.71
-Mesonet 14.8 6.67 28.57 22.86 54.76 36.18 44.29 49.81 46.94 33.88
-Xception 1.71 0 9.52 17.14 22.21 28.82 9.29 40.75 29.16 17.62
-MultiAtten 4.13 0 28.57 13.81 44.42 30.32 35.96 43.93 43.57 27.19
-CDCN++ 0.28 6.67 10.71 31.9 43.83 33.19 8.39 44.58 29.68 23.25
-DeepPixel 0.14 0 10.12 13.33 40.19 32.8 2.14 37.38 19.48 17.29
-ResNet50 1.21 0 6.55 12.86 28.79 21.55 4.29 36.82 23.06 15.01
-ViT 0 0 0 9.52 46.3 16.2 2.68 30.99 20.01 13.97
-JFSFDB-Uni 0 0 0.6 6.67 43.6 11.64 2.32 30.6 18.8 12.69
-CLIP-V 0 0 10.71 8.33 37.17 23.46 11.43 25.72 27.81 16.8
-CLIP 0 0 8.13 10.24 23.58 26.86 12.14 23.76 27.34 15.4
-CoOp(Unified) 0.02 0 11.11 28.57 17.77 14.39 28.57 36.69 39.84 19.66
-CoOp(Specific) 1.54 0 11.11 29.76 13.82 14.48 31.43 37.22 40.31 19.96
-EER(%) Joint HipTune 0 0 0 3.84 11.72 8.99 1.23 24.92 12.21 6.99
-TABLE X: Results on JFSFDB-EER.
-JFSFDB primarily demonstrated that training with a combination of physical and digital attack samples can significantly enhance model performance.
-This conclusion is also applicable to our HiPTune model.
-When trained and tested separately within the dataset, HiPTune achieved an EER of 9.71% and an AUC of 93.12%.
-However, after joint training, HiPTune gained access to a more diverse set of classification criteria, which it could adaptively select and integrate, resulting in a substantial performance improvement: a 4.34% increase in AUC and a 2.72% reduction in EER.
-Moreover, the model’s generalization capability was significantly enhanced.
-V-E Results on UniAttackData
-Protocol Methods ACER ACC AUC EER Protocol Methods ACER ACC AUC EER
-ResNet-50 1.35 98.83 99.79 1.18 ResNet-50 34.6 53.69 87.89 19.48
-ViT-B/16 5.92 92.29 97 9.14 ViT-B/16 33.69 52.43 83.77 25.94
-Auxiliary 1.13 98.68 99.82 1.23 Auxiliary 42.98 37.71 76.27 32.66
-CDCN 1.4 98.57 99.52 1.42 CDCN 34.33 53.1 77.46 29.17
-FFD 2.01 97.97 99.57 2.01 FFD 44.2 40.43 80.97 26.18
-CLIP-V 0.46 99.55 99.95 0.46 CLIP-V 11.38 85.93 95.73 11.38
-CLIP 0.58 97.4 99.93 0.58 CLIP 15.27 79.83 92.46 15.27
-CoOp(unified) 7.94 90.96 97.9 7.94 CoOp(unified) 33.29 63.89 71.58 33.29
-CoOp(specific) 8.49 90.98 97.45 8.49 CoOp(specific) 35.67 60.9 68.41 35.67
-UniAttackDetection 0.52 99.45 99.96 0.53 UniAttackDetection 22.42 67.35 91.97 15.72
-1 HiPTune 0.33 99.98 99.99 0.36 2 HiPTune 10.38 88.62 97.82 10.39
-TABLE XI: Results on UniAttackData.
-In the UniAttackData dataset, where the data distribution is relatively simple, our HiPTune model achieved outstanding performance.
-Leveraging the guidance function of prompts, HiPTune also delivered excellent results in cross-attack protocols, demonstrating its robustness and adaptability across different attack types.
-V-F Abalation Study
-To verify the effectiveness of the theory proposed by HiPTune and to analyze the influence of various settings, we conduct several ablation studies on protocol $P_{2}$ of HiFAB.
-These studies examine the depth of the tree, length of prompts, different selection strategies, various gate functions, and the contribution of each component.
-Additional experiments and detailed analyses of each element are provided in the Supplementary Materials.
-Depth of Tree.
-The depth of the VP-Tree plays a crucial role in helping models understand fine-grained classification criteria, which enhances their ability to handle unseen advanced attacks during the integration process.
-Models such as CoOp, which lack a hierarchical structure, exhibit poor performance as they rely on a single, rigid classification criterion.
-In contrast, increasing the depth of the VP-Tree allows the model to ”draw” the line between live and fake faces adaptively and accurately, demonstrating the importance of a hierarchical approach.
-Length of Prompts.
-The length of prompts determines the scope of learnable parameters, impacting both accuracy and efficiency.
-To explore this balance, we conduct several experiments to assess the effect of different prompt lengths.
-Using more parameters can capture more detailed classification-related information, enhancing model performance.
-Component ACER(%) AUC(%) EER(%) ACC(%)
-level-0 (CoOp) 14.20 92.57 14.26 85.44
-level-1 9.78 94,26 10.12 90.12
-level-2 5.78 96.88 5.45 95.08
-level-3 3.26 98.78 3.67 95.79
-TABLE XII: The influence of depth of the VP-Tree.
-Component ACER(%) AUC(%) EER(%) ACC(%)
-16 4.24 95.79 4.26 92.39
-20 4.08 96.03 4.03 93.40
-29 3.45 96.34 3.44 94.67
-40 3.48 97.62 3.90 94.23
-60 3.26 98.78 3.67 95.79
-TABLE XIII: The impact on the length of prompts.
-VI Conclusion
+The initialization process of VP-Tree has been trained for $100$ epochs (batch size: $32$), and followed by full model training for $500$ epochs (batch size: $32$).
+The image tokens are first fed into $N_{0}=3$ blocks to gain the initial representation abilities.
+In our case, $N_{1}=N_{2}=N_{3}=3$ blocks are followed by each selection of prompts.
+TABLE V: Results on the JFSFDB with three cross-domain tasks: PAD, FFD, and UAD.
+For PAD, the training set consists of SiW, 3DMAD, and HKBU, while the testing set includes MSU, 3DMask, and ROSE.
+For FFD, the training set is FF++, and the testing set comprises DFDC, CelebDFv2.
+For UAD, the training set combines SiW, 3DMAD, HKBU, and FF++, and the testing set includes MSU, 3DMask, ROSE, DFDC, and CelebDFv2.
+↓/↑ indicates that smaller/larger values correspond to better performance.
+The Avg.
+represents the mean result for all testing sets.
+Metrics Method PAD FFD UAD Avg.
+MSU 3DMask ROSE DFDC Celeb-DFv2 MSU 3DMask ROSE DFDC Celeb-DFv2
+EER(%)↓ MesoNet [93] 21.90 55.82 36.81 46.16 38.78 22.86 54.76 36.18 49.81 46.94 41.00
+Xception [94] 18.57 42.30 18.13 38.92 24.07 17.14 22.21 28.82 40.75 29.16 28.00
+CDCN++ [89] 22.80 52.76 32.01 44.75 28.19 31.90 43.83 33.19 44.58 29.68 36.37
+DeepPixel [6] 15.71 46.42 28.23 36.82 22.81 13.33 40.19 32.80 37.38 19.48 29.32
+JFSFDB-Uni [38] - - - - - 6.67 43.60 11.64 30.60 18.80 22.26
+CLIP [83] 12.86 26.74 14.61 23.14 27.16 10.24 23.58 26.86 23.76 27.34 21.63
+CoOp [86] 25.95 17.02 12.06 39.42 41.21 28.57 17.77 14.39 36.69 39.84 27.29
+FA3-CLIP [59] 6.90 18.87 9.61 22.46 25.51 4.29 24.00 10.77 24.53 30.02 17.70
+HipTune 6.56 11.57 12.82 20.86 28.77 3.84 11.72 8.99 24.92 12.21 14.23
+AUC(%)↑ MesoNet [93] 85.33 44.54 68.05 54.72 65.10 84.01 45.87 70.00 50.57 55.03 62.32
+Xception [94] 90.62 63.88 90.15 64.86 84.27 91.56 75.49 79.38 63.69 78.87 78.28
+CDCN++ [89] 82.65 47.66 76.76 56.65 78.34 78.93 61.54 76.39 56.42 76.92 69.23
+DeepPixel [6] 95.17 56.87 80.60 67.54 85.51 93.97 66.65 74.58 66.36 88.45 77.57
+JFSFDB-Uni [38] - - - - - 97.99 67.10 95.57 75.90 89.76 85.26
+CLIP [83] 95.82 80.40 92.89 86.08 79.40 96.39 86.13 81.07 84.06 78.99 86.12
+CoOp [86] 83.73 91.94 94.63 64.46 62.05 81.58 89.27 93.62 68.56 64.88 79.47
+FA3-CLIP [59] 97.96 89.99 95.52 85.60 83.15 98.12 81.84 95.58 82.95 74.80 88.55
+HipTune 98.62 95.87 98.62 89.72 89.63 99.62 95.43 97.82 90.88 93.35 94.96
+5.2 Results on UniAttackData+
+We employ a range of CLIP-based baselines, including CLIP [83], CLIP-V [83], and CoOp [86] to evaluate the performance in UAD on our dataset.
+In addition, we assess algorithms specifically designed for PAD/FFD, such as CFPL-FAS [90], MoE-FFD [91], and LLA-Net [92], to examine their generalization capabilities.
+We also compare our approach with the recently released FA3-CLIP, an algorithm tailored for unified attack detection.
+The detailed experimental results are presented in Tab.
+IV.
+Protocol 1 is designed to evaluate the model’s generalization of face identity in the UAD task.
+As shown in Tab.
+IV, we can observe that CLIP-V [83] achieves decent performance, with an ACER of 9.36%.
+However, introducing fixed text templates or single-semantic learnable vectors does not bring performance benefits, such as the ACER of CLIP [83] and CoOp [86] being 9.93% and 14.76%, respectively.
+These experimental results indicate that in the UAD task, naive category prompts cannot accurately describe highly challenging multi-attack types, resulting in ineffective differentiation from real faces.
+Although the recently released CDCN++[89], CFPL-FAS[90], MoE-FFD [91], and LLA-Net [92] have achieved superior performance in PAD and FFD tasks, their effectiveness drops significantly on our UniAttackData+ dataset, with ACERs of 12.77%, 13.93%, 17.30%, and 10.57%, respectively.
+This suggests that these algorithms exhibit poor generalization across different attack categories, as models trained on PAD datasets fail to adapt effectively to FFD scenarios.
+FA3-CLIP [59] solves UAD by combining spatial and frequency features in a dual-stream architecture.
+By employing an attack-agnostic prompt learning and bridging the gap between different attack types, it significantly improves UAD performance, such as ACER is 8.26%.
+Our proposed HiPTune employs a visual prompt-based strategy that effectively guides the encoder toward informative features while preserving model efficiency.
+Furthermore, the hierarchical prompt tree captures multi-level, fine-grained classification criteria, enabling highly accurate and generalizable performance across diverse attack scenarios, with an ACER of just 7.31%.
+Protocol 2 evaluates the generalization of the model across attack methods in the UAD task, which is widely applicable in real-world scenarios.
+When encountering unseen attacks, algorithms that either lack text prompts or rely solely on simple prompts exhibit significantly degraded performance.
+For example, CLIP-V [83] and CoOp [86] achieve ACERs of 28.77% and 43.14%, respectively.
+Although FA3-CLIP [59] demonstrates promising results on Protocol 1, its effectiveness significantly declines when encountering unknown attack types, with an ACER of 35.81%.
+This suggests that single-semantic prompt learning struggles to generalize across diverse attack categories.
+Unlike conventional prompt-matching methods, our prompts function solely as guidance rather than decision-making anchors, thereby reducing reliance on prior knowledge.
+This design enhances the model’s generalization to unseen attack methods while ensuring robustness and interpretability, achieving an ACER of 14.26%.
+Protocol 3 is designed to evaluate whether models trained on simple attack types can retain high robustness when tested against complex and sophisticated attacks.
+We can observe that almost all methods have degraded performance, even to the point of random guessing, such as the ACER of CLIP-V [83], CoOp [86], MoE-FFD [91], and FA3-CLIP [59] exceeding 40%.
+In contrast, our HiPTune model adaptively integrates diverse types of prompts and demonstrates superior performance.
+Despite being trained on limited samples, HiPTune leverages high-level prompts and auxiliary information from other attack types to support inference.
+This design enables the model to maintain strong generalization capabilities, even against advanced attacks, achieving an ACER of 25.68%.
+TABLE VI: Results on UniAttackData with standard Protocol 1, 1.1, 1.2, 1.3, and 2.
+↓/↑ indicate that smaller/larger values correspond to better performance.
+The Avg.
+represents the mean ACER for Protocol 1, 1.1, 1.2, 1.3.
+Methods P1 P1.1 P1.2 P1.3 Avg. P2
+ACER(%)↓ AUC(%)↑ ACER AUC ACER AUC ACER AUC ACER AUC
+ResNet-50 1.35 99.79 5.92 91.25 25.90 84.35 4.92 98.84 9.52 34.60 87.89
+ViT-B/16 5.92 97.00 13.53 95.99 5.22 99.36 3.20 99.18 6.97 33.69 83.77
+CDCN [89] 1.40 99.52 12.32 93.89 16.34 93.34 4.41 97.68 8.62 34.33 77.46
+CLIP [83] 1.02 99.47 14.81 86.74 5.36 99.17 2.45 97.92 5.91 15.27 92.46
+UniAttackDetection [39] 0.52 99.96 11.73 98.81 1.70 99.85 4.67 99.13 4.66 22.42 91.97
+FA3-CLIP [59] 0.36 99.75 9.57 97.78 1.43 99.85 2.30 99.19 3.42 18.81 88.59
+HiPTune 0.33 99.99 7.63 90.52 1.26 98.88 1.08 97.18 2.58 10.38 97.82
+5.3 Results on JFSFDB Dataset
+The Tab.
+V presents the evaluation results of various methods across three tasks: PAD, FFD, and UAD.
+Two metrics are used for performance comparison: Equal Error Rate (EER%) and Area Under Curve (AUC%).
+The experiment compares traditional methods (e.g., MesoNet [93], Xception [94], DeepPixel [6], CDCN++ [89], JFSFDB-Uni [38]), CLIP-based baselines (CLIP [83], CoOp [86], FA3-CLIP [59]), and the proposed method HiPTune.
+In tasks PAD and FFD, our HiPTune achieves the best results in the AUC metric, such as 98.62%, 95.87%, and 98.62% on datasets MSU, 3DMask, and ROSE, and 89.72% and 89.63% on datasets DFDC and Celeb-DFv2, respectively.
+However, in terms of EER, HiPTune shows slightly weaker generalization on the ROSE and Celeb-DFv2 datasets.
+We attribute this to a class imbalance issue, where the model tends to learn better representations for the majority class (e.g., real faces).
+As a result, it achieves strong overall ranking performance (reflected in a high AUC).
+However, during testing, the decision threshold cannot effectively balance the trade-off between the two classes, leading to suboptimal EER performance.
+After training the UAD model with joint physics and digital forgery data, our HiPTune further improved the performance of AUC on most cross-dataset test results, such as MSU improving by 1.00% (98.62%→99.62%), DFDC improving by 1.16% (89.72%→90.88%), and Celeb-DFv2 improving by 3.72% (89.63%→93.35%).
+These experimental results indicate that our HiPTune can combine the diversity and complementarity of PAs and DFs to improve its performance in PAD and FFD tasks, respectively.
+However, for datasets 3DMask and ROSE, the performance of UAD slightly decreased, such as 95.87% decreasing to 95.43%, 98.62% decreasing to 97.82%, respectively.
+This indicates that introducing the FF++ dataset into the training set can impair the model’s ability to recognize 3D masks.
+We analyze that in our HiPTune framework, prompts are organized hierarchically by attack type.
+If a large influx of digital attacks (like from FF++) dominates the prompt learning phase, it could bias the learned prompts or pruning decisions, weakening the semantic alignment for physical branches (like 3D masks).
+5.4 Results on UniAttackData
+Tab.
+VI presents the performance of various methods on the UniAttackData under standard Protocols 1, 1.1, 1.2, 1.3, and 2, where lower ACER and higher AUC indicate better results.
+The Avg.
+column summarizes the average ACER across Protocols 1, 1.1, 1.2, and 1.3.
+HiPTune achieves the best overall performance, with the lowest average ACER (2.58%) and the highest AUC (99.99%) under Protocol 1, as well as strong performance in Protocol 2 (ACER: 10.38%, AUC: 97.82%).
+On Protocol 1, 1.2, and 1.3, HiPTune achieves the best ACER values of 0.33%, 1.26%, and 1.08%, respectively, and obtains the best mean ACER (2.58%).
+It indicates that HiPTune has the best generalization for facial identity information, such as the strict identity partitioning in the training, validation, and testing sets of Protocol 1.3.
+When physical presentation attacks and adversarial attacks are used as training sets, HiPTune has poorer generalization to deepfake attacks than ResNet-50, such as 7.63% v.s.
+5.92% for ACER.
+On the contrary, the performance of ResNet-50 has significantly decreased, such as achieving an ACER value of 25.90% on Protocol 1.2.
+However, HiPTune still maintains its leading performance, achieving an ACER of 1.26%.
+This proves the effectiveness of using hierarchical prompt learning in HiPTune, where the model uses a hierarchical routing mechanism to select the most relevant prompts for each input image.
+Finally, on the cross-attack category Protocol 2, where the model is trained on physical presentation attacks and tested on digital forgery (including opposite training and testing), our HiPTune demonstrates an absolute advantage, achieving an ACER of 10.38% and an AUC of 97.82%.
+HiPTune demonstrates strong generalization across different attack categories, due to the following points: (1) Hierarchical prompt modeling captures shared and distinct semantics.
+VP-Tree organizes attacks hierarchically by semantic similarity.
+High-level prompts (“physical” vs.“digital”) provide coarse but generalizable decision boundaries, while low-level prompts capture fine-grained modality-specific traits.
+This structure allows the model to transfer knowledge from one modality (e.g., physical) to another (e.g., digital), by reusing the shared higher-level decision boundaries.
+(2) Dynamic prompt selection tailored to each sample.
+The APP module ensures that, at inference time, the model can select the most appropriate semantic pathway, even if that pathway wasn’t explicitly optimized during training.
+This enables robust out-of-distribution generalization when tested on unseen attack types.
+5.5 Results on cross-domain UAD benchmark
+TABLE VII: Results on cross-domain UAD benchmark.
+Method U&U+→J J&U+→U J&U→U+
+ACER AUC ACER AUC ACER AUC
+CLIP-V 45.12 56.26 13.15 91.86 39.81 63.4
+CLIP 44.42 58.10 16.59 90.14 39.36 64.08
+CoOp 38.21 65.84 30.28 76.11 29.77 76.77
+HiPTune 32.93 70.89 8.55 94.67 25.93 83.05
+Tab.
+VII presents the results of various methods, i.e., CLIP-V [83], CLIP [83], CoOp [86], and HiPTune on the cross-domain UAD benchmark, designed to evaluate the generalization ability of the models under domain shifts.
+The benchmark consists of three protocols, each involving training and testing across different combinations of datasets: UniAttackData (U), UniAttackData+ (U+), and JFSFDB (J).
+Among the compared methods, HiPTune achieves the best performance across all three cross-domain settings.
+For example, when the test set is J, U, and U+, our HiPTune achieves ACER values of 32.93%, 8.55%, and 25.93%, respectively, which are significantly better than the compared algorithms.
+The superiority of domain generalization in our analysis comes from the following two points: (1) HiPTune constructs a VP-Tree that models attack types in a coarse-to-fine hierarchy, enabling the model to distinguish diverse attack semantics across different domains.
+This structured representation ensures that even if an unseen domain presents a new variant of an attack, the model can still leverage higher-level prompts to make correct predictions.
+(2) The APP module dynamically selects the most appropriate prompts for each sample based on its visual features, rather than relying on fixed global parameters.
+This flexibility allows the model to adapt to distribution shifts without retraining, thereby improving generalization to unseen domains.
+5.6 Ablation Studies
+Effect of VP-Tree Depth.
+To evaluate the contribution of the hierarchical structure in VP-Tree, we conduct experiments on protocol 1 of UniAttackData+ by varying the depth of the tree from 1 to 3.
+As shown in Fig.
+5, deeper trees consistently improve the performance across all metrics.
+Specifically, the full three-level VP-Tree achieves the best ACER (7.31%) and EER (7.61%), indicating that more fine-grained semantic grouping leads to more robust classification criteria.
+This demonstrates the benefit of modeling attack diversity hierarchically.
+Effect of Prompt Length.
+We also study the impact of the prompt length in our framework.
+As presented in Fig.
+5, increasing the length from 10 to 40 improves performance steadily.
+The model achieves optimal results at a length of 40, with an ACER of 7.31% and an AUC of 97.87%.
+Further increasing the length to 50 offers marginal gains in accuracy but slightly degrades ACER and EER, suggesting a trade-off between prompt expressiveness and overfitting.
+Thus, we adopt 40 as the default length in our experiments.
+Figure 5: Ablation studies on the level of VP-Tree and the length of the prompt.
+6 Conclusion
 In this paper, we first present a large-scale Unified Face Attack Benchmark (HiFAB), which is the most extensive and comprehensive UAD dataset to date.
@@
 Every attack is applied to $2,875$ identities to reduce the ID noise.
-Furthermore, we introduce a novel Hierarchical Prompt Tuning framework (HiPTune), that adaptively integrates classification criteria from various semantic spaces, allowing tailored classification criteria for each sample.
-Finally, we conduct several experiments that evaluate the model’s performance in real-world scenarios.
-VII Disscussion and Future Work
-In this work, we explore the application of prompt tuning in the visual domain while acknowledging the value of examining its adaptability to its original text-based purpose.
-The hierarchical and semantic capabilities of text-based prompt tuning could provide valuable insights for visual applications, warranting further investigation.
-Furthermore, the suboptimal performance of baseline models under protocol $P_{1}$ may be partially attributed to the absence of a fine-tuning process on the testing datasets.
-This lack of adaptation limits the models’ ability to generalize to unseen attack types and domain variations.
-Further discussion on the impact of fine-tuning in bridging this gap will be presented.
+Furthermore, we introduce a novel Hierarchical Prompt Tuning framework (HiPTune) that adaptively integrates classification criteria from various semantic spaces, allowing tailored classification criteria for each sample.
+Finally, the proposed HiPTune demonstrates superior performance and generalization across three unified face forgery detection benchmarks.